### v1

# LLMs Cannot Reliably Judge (Yet?): A Comprehensive Assessment on the Robustness of LLM-as-a-Judge

Abstract
Large Language Models (LLMs) have demonstrated remarkable intelligence across various tasks, which has inspired the development and widespread adoption of LLM-as-a-Judge systems for automated model testing, such as red teaming and benchmarking. However, these systems are susceptible to adversarial attacks that can manipulate evaluation outcomes, raising concerns about their robustness, hence trustworthiness. Existing evaluation methods adopted by LLM-based judges are often piecemeal and lack a unified framework for comprehensive assessment. Furthermore, the prompt template and model selections for improving judge robustness have been rarely explored, and their performance in real-world settings remains largely unverified. To address these gaps, we introduce RobustJudge, a fully automated and scalable framework designed to systematically evaluate the robustness of LLM-as-a-Judge systems. RobustJudge investigates the impact of attack methods and defense strategies (RQ1), explores the influence of prompt template and model selection (RQ2), and assesses the robustness of real-world LLM-as-a-Judge applications (RQ3). Our main findings are that 1) LLM-as-a-Judge systems are still vulnerable to a range of adversarial attacks, including Combined Attack and PAIR, while defense mechanisms such as Re-tokenization and LLM-based Detectors offer improved protection. 2) Robustness is highly sensitive to the choice of prompt template and judge models. Our proposed prompt template optimization method can improve robustness, and JudgeLM-13B demonstrates strong performance as a robust open-source judge. 3) Applying RobustJudge to Alibaba’s PAI platform reveals previously unreported vulnerabilities. The source code of RobustJudge is provided at https://github.com/S3IC-Lab/RobustJudge.
1 Introduction
Large Language Models (LLMs), such as OpenAI’s GPT-4o [1], Google’s Gemma2 [2], Meta’s Llama 3 [3], and QwenLM’s Qwen2.5 [4], have achieved remarkable proficiency across a wide range of tasks. Built on extensive training data and Transformer architectures, these models demonstrate advanced capabilities in general-purpose intelligence, including natural language understanding, text generation, and complex problem-solving.
To leverage these capabilities while minimizing human effort and mitigating human biases in LLM evaluation, the concept of LLM-as-a-Judge has been introduced. LLM-as-a-Judge [5] aims to automate the assessment of LLM-generated content, providing a scalable and objective alternative to human evaluation. This approach has gained widespread adoption, becoming a defacto evaluation method for assessing LLM performance across various domains, including software engineering[6], domain-specific knowledge assessment[7], and mathematical reasoning[8].
The initial success of LLM-as-a-Judge can be attributed to its strong agreement with human preferences on standard benchmarks. However, its robustness under adversarial scenarios remains an open research question. Recent studies [9, 10, 11, 12] have revealed that LLM-as-a-Judge systems are intrinsically vulnerable to various forms of adversarial attacks, which can manipulate evaluation outcomes without introducing easily detectable anomalies. These findings raise significant concerns about the reliability of LLM-as-a-Judge systems.
Given the increasing reliance on LLM-as-a-Judge platforms for LLM evaluation, such as AlpacaEval [13], Chatbot Arena [5], and MT-Bench [5], ensuring their robustness has become a pressing research priority. However, existing assessment methods for these systems are fragmented, lacking a unified, systematic, and automated framework for comprehensive evaluations. Furthermore, the optimal prompt configuration and model selection for different tasks and evaluation protocols are underexplored, with limited guidance on best practices. Finally, while real-world LLM-as-a-Judge applications have recently demonstrated promising effectiveness in content evaluation, their robustness in adversarial settings remains unverified.
Our Work. To address these challenges, we introduce RobustJudge, a fully automated and scalable framework to evaluate the robustness of LLM-as-a-Judge systems. Our framework systematically assesses these systems by exploring three core research questions:
  • RQ1: What impact do different adversarial attacks and defense methods have on the LLM-based judges?
  • RQ2: How do the prompt templates and model choices affect the robustness of LLM-based judges?
  • RQ3: What vulnerabilities exist in black-box real-world deployments of LLM-based judges, as revealed by our empirical evaluations?
To answer RQ1, we conduct a comprehensive evaluation of LLM-based judges against 15 adversarial attack techniques and 7 defense strategies. Our analysis provides extensive comparisons of these techniques, revealing critical insights into their relative strengths and weaknesses. The results indicate that many attack methods, such as Fake Reasoning [12], Combined Attack [14], Empty Attack and AdvEval [10], PAIR [15], TAP [16] consistently achieve high attack success rate on multiple tasks and models, highlighting vulnerabilities in LLM-based judges. Conversely, defense strategies such as Retokenization [17] and Naive LLM-based Detector [18], demonstrate notable effectiveness in mitigating these attacks. Our evaluation provides a comprehensive view of the current landscape of attack and defense techniques, and a clear guidance on safeguarding LLM-as-a-Judge systems against adversarial manipulation.
To address RQ2, we analyze the effect of adversarial attacks under different judge prompt templates and judge model selection. Our analysis reveals that the robustness of LLM-as-a-Judge systems is highly sensitive to both factors. Specifically, while all the evaluated prompt templates are clear and well-structured, they present different levels of robustness to the LLM-as-a-Judge system. To mitigate this issue, we propose a prompt template optimization method aimed at identifying configurations with improved robustness. Our optimized template consistently outperforms existing templates in robustness against multiple attacks. Additionally, our evaluation on various judge models reveals that JudgeLM-13B exhibits substantially stronger robustness against adversarial inputs, performing comparably to the widely adopted GPT-4o. This suggests that JudgeLM-13B is a strong open-source alternative as an LLM judge. These findings emphasize the importance of prompt template design and judgment fine-tuning in enhancing the robustness of LLM-as-a-Judge systems against adversarial attacks.
To study RQ3, we investigate the robustness of a real-world LLM-based judge system, i.e., Alibaba’s PAI platform. Our evaluation leverages the adversarial examples generated by RobustJudge and is conducted via the public API provided by the platform. The results show that conventional adversarial attacks were ineffective against PAI-Judge platform, suggesting a strong level of robustness. However, we identify a critical loophole using a composite attack, which combines PAIR-optimized adversarial inputs with long-suffix manipulations. This strategy successfully bypasses the platform’s defenses and alters the model’s evaluation outcome. These findings indicate that RobustJudge is practical in identifying hidden vulnerabilities and guiding the development of more robust LLM-as-a-Judge systems.
This paper makes the following key contributions:
  • We develop RobustJudge, the first fully automated and scalable framework designed for extensive robustness evaluation of LLM-as-a-Judge systems.
  • RobustJudge evaluate 15 adversarial attacks and validates 7 judge defense strategies. Our analysis reveals vulnerabilities of LLM judges on various tasks, consistency of judgment across evaluation protocols, and high attack success rates from methods such as Combined Attack and PAIR. While defenses like Re-tokenization and LLM-based Detectors show promise, their effectiveness often comes with notable trade-offs in accuracy or usability.
  • Utilizing RobustJudge, we conduct an in-depth investigation on the configuration of LLM-as-a-Judge systems, focusing on the prompt templates and judge model selection. Our extensive analysis identifies the most reliable configuration against adversarial attacks, offering actionable guidance for enhancing system robustness.
  • We evaluate a real-word LLM judge system Alibaba’s PAI platform using RobustJudge, and identify hidden vulnerability of the platform. We share our findings with the PAI team to support the development of more robust LLM-as-a-Judge systems.
2 Background and Related Work
2.1 LLM-as-a-Judge
Large language models (LLMs) have exhibited state-of-the-art proficiency in understanding and generating human-like text. Despite their impressive capabilities, traditional reference-based metrics such as ROUGE [19] and BLEU [20] are limited in capturing the nuanced quality of their outputs. To overcome these limitations, Zheng et al. proposed the LLM-as-a-Judge paradigm [5], which employs LLMs themselves as evaluators to assess the quality of model outputs. This paradigm has rapidly gained popularity, leading to the development of a variety of fine-tuned judge models tailored for different evaluation settings [21].
For example, Wang et al. developed PandaLM [22], which avoids costly API calls and mitigates privacy risks. Zhu et al. proposed JudgeLM [23], enhancing accuracy via swap augmentation and reference support. Kim et al. released Prometheus 2 [24], an open-source evaluator achieving top alignment with human and GPT-4 judgments. Beyond general text, the judge paradigm has been adapted to specialized domains, including software engineering [6], machine translation [25], legal reasoning [7], and mathematical problem solving [8]. Notably, Zhuge et al. introduced Agent-as-a-Judge [26], leveraging autonomous agents as evaluators.
Several benchmarks have also been developed to evaluate the judge’s performance. Early efforts like LLMEval [27], MTBench [5], and FairEval [28] focus on consistency with human judgment, though they may favor stylistic fluency over factual correctness. LLMBar [29] improves evaluation rigor by introducing ground-truth preference labels and stricter instruction adherence. JudgeBench [8] further advances the field by emphasizing objective correctness in challenging domains such as knowledge, reasoning, mathematics, and programming.
However, while these benchmarks assess overall judging ability, they do not examine evaluator safety or robustness. In this work, we introduce the first benchmark designed specifically to evaluate the safety risks and reliability of LLM-based evaluators.
Figure 1: LLM-as-a-judge pipeline.
2.2 Attacks against LLM-as-a-Judge
We categorize attacks on LLM-as-a-Judge into two main types: heuristic-based and optimization-based.
Heuristic-based Attacks. Heuristic-based attacks refer to adversarial strategies that are manually crafted or rule-based, exploiting model behaviors without relying on gradient information or automated optimization. Typically, these attacks leverage prompt engineering and input manipulation to coerce the model into unintended behaviors, and can be broadly categorized into prompt injection attacks and adversarial example attacks.
In prompt injection attacks, a basic approach is to concatenate the benign input with an adversarial sequence [30, 31, 32]. Alternatively, inserting special tokens or delimiters can activate hidden or unauthorized behaviors [31].
Optimization-based Attacks. Optimization-based attacks leverage gradient signals or structured search procedures to systematically craft inputs that achieve specific adversarial objectives. In this section, we focus on both jailbreak attacks and adversarial strategies.
Beyond prompt injection, Chen et al. [10] introduced AdvEval, a black-box adversarial framework that generates targeted examples against NLG evaluators using feedback from powerful LLMs. By iteratively refining inputs based on discrepancies between human and model judgments, AdvEval substantially degrades the performance of multiple evaluation metrics across diverse tasks and datasets.
Jailbreaking involves designing inputs that circumvent an LLM’s built-in safeguards, coercing it into producing content that violates its usual policy constraints, such as harmful or unethical outputs [33, 16, 34, 15]. However, conventional jailbreak techniques cannot be directly applied to LLM-as-a-Judge, since the evaluation setting does not reduce to simple suffix optimization. To bridge this gap, we propose a novel adaptation framework that integrates existing jailbreak methodologies into the LLM-as-a-Judge context, enabling targeted manipulation of the judge’s scoring behavior.
Recent studies have directly targeted LLM-based evaluators via adversarial manipulations. Zheng et al. [11] demonstrated that a null model—always outputting a constant, irrelevant response—can still secure high win rates on benchmarks like AlpacaEval 2.0 and MT-Bench, exposing fragility of benchmark-driven evaluation. Raina et al. [9] revealed that appending short, task-agnostic adversarial phrases can dramatically inflate scores in absolute scoring setups, underscoring the lack of robustness in LLM judges. Shi et al. [12] introduced JudgeDeceiver, a gradient-based prompt injection method that optimizes adversarial sequences to reliably mislead LLM evaluators across tasks—ranging from LLM-powered search to RLAIF—surpassing the effectiveness of manual prompt attacks.
In this paper, we integrate a wide range of adversarial strategies into a unified framework to systematically evaluate their effectiveness against LLM-as-a-Judge systems.
2.3 Defenses for LLM-as-a-Judge
We categorize defense strategies against attacks on LLM-as-a-Judge into two main types: prevention-based and detection-based approaches.
Prevention-based Defenses. Prevention-based defenses seek to block prompt injection and adversarial manipulation by preprocessing both input instructions and candidate outputs before evaluation. For instruction preprocessing, methods have been developed to refocus the model on the genuine task and neutralize injected prompts. For instance, Sandwich [35] inserts an instruction-reinforcing prompt after the input to maintain task integrity. For candidate response processing, Jain et al. [17] employed paraphrasing and retokenization to obscure adversarial triggers, and Liu et al. [14] further adapted these techniques to counter more sophisticated injections. Additionally, Li et al. [36] used LLMs to mask and reconstruct the input, disrupting adversarial continuity and substantially lowering attack success rates. Nevertheless, these generation-focused strategies often prove inadequate for LLM-as-a-Judge, where attackers target score distortions rather than harmful content generation.
Detection-based Defenses. Detection-based defenses focus on identifying potential attacks at either the input or output stages of evaluation. On the input side, Jain et al. [17] introduced a self-perplexity filter to detect prompt injection by identifying anomalous perplexity signals. Follow-up work [37] enhanced this approach by combining perplexity and token-length features to train a classifier for detecting adversarial prompts. While these input-level methods offer some protection, they often struggle to generalize to more sophisticated adversarial strategies—such as optimization-based prompt injections targeting judge LLMs [12] or adversarial input crafting techniques [9, 10] that directly manipulate score assignment or response ranking.
On the output side, Helbing et al. [38] proposed using an auxiliary LLM to verify whether generated outputs exhibit jailbreak behaviors. Similarly, sandwich-based detection [35] evaluates whether outputs remain aligned with predefined system objectives. However, such output-level defenses are generally less effective in the judge setting, where both benign and adversarial candidate responses may appear plausible. In these cases, the attack’s goal is not to generate visibly harmful content but to subtly distort scoring or ranking decisions.
In this paper, we extend existing defenses to a broad range of LLM-as-a-Judge tasks and evaluate their robustness against adversarial attacks.
3 Threat Model
We identify two principal roles: the attacker, who crafts attacks against LLM-as-a-Judge, and the defender, who safeguards the evaluation’s reliability and safety.
3.1 Attacker
We categorize adversaries against LLM-as-a-Judge systems into two classes: heuristic-based and optimization-based attackers. Heuristic-based adversaries, lacking direct access to the evaluation metric $\mathit{m}$, craft adversarial inputs using indirect cues, such as score feedback [10] or prompt manipulation techniques [31, 30, 32, 39, 40, 14]. In contrast, optimization-based adversaries have full visibility into the evaluation prompt and scoring rules, enabling white-box optimization to directly maximize their evaluation outcome [11, 12, 9].
Attacker’s Capabilities. Regardless of attacker type, we assume the attacker can manipulate the response submitted the LLM judge for evaluation. In particular, the attacker is allowed to inject arbitrary instructions or sequences into the malicious response. For example, the attacker may insert additional content at any position within the response to increase its score or the likelihood of being selected. However, attackers are not allowed to modify the judge’s instruction prompt or internal model parameters, which are fixed and controlled by the host institution. Besides, attackers do not have access to the competing candidate’s response $r$ in pairwise evaluation settings.
Attacker’s Goal. The attacker’s goal is to manipulate the LLM-as-a-Judge pipeline by crafting responses that skew evaluations in their favor.
In the scoring scenario, the attacker generates a malicious response $r_{m}$ designed to artificially maximize the probability of the top score without genuinely meeting the criteria. Formally, it solves:
$\arg\max_{w_{1},\dots,w_{k}\in V_{\text{score}}^{k}}P(w_{1},\dots,w_{k}\mid\mathcal{E}(q,m,r_{m}))=w_{k},$
where $w_{k}$ is the token for the highest evaluation score.
In the pairwise comparison scenario, the attacker forges $r_{m}$ to ensure it is chosen over a benign response $r_{b}$, i.e.:
$\arg\max_{w\in V_{\text{pair}}}P(w\mid\mathcal{E}(q,m,r_{m},r_{b}))=w_{m},$
where $w_{m}$ indicates selection of the malicious response.
3.2 Defender
We categorize defenses for LLM-as-a-Judge into two main types: prevention and detection. Prevention-based defenses aim to redesign the evaluation prompt or pre-process inputs to block adversarial manipulations from influencing the model’s judgments. Detection-based defenses focus on identifying whether the evaluation data has been tampered with or compromised.
Defender’s Capability. The defender cannot control the attacker’s actions or modify the model post-release. In closed-source settings, the defender may additionally implement dynamic defense mechanisms, including real-time detection of compromised inputs, instruction reinforcement to mitigate prompt injection attempts, and adaptive intervention during evaluation. Across both settings, the defender assumes that attackers can freely submit malicious candidate responses but cannot modify the judge’s internal model weights or core prompt instructions.
Defender’s Goal. The defender’s goal is to safeguard the evaluation integrity of LLM-as-a-Judge systems by ensuring that judgments are based solely on the intended evaluation criteria, without being influenced by adversarial manipulations. Specifically, the defender aims to (i) prevent the execution of injected instructions through prompt design and input sanitization, and (ii) detect and mitigate compromised evaluation data when prevention fails.
An effective defense should preserve evaluation fidelity under both benign and adversarial conditions while minimizing utility degradation and false detections.
4 RobustJudge
4.1 Overview of RobustJudge
Figure 2: Overview of the RobustJudge workflow. RobustJudge is a framework to evaluate the robustness of LLM-as-a-Judge systems. It supports customizable configuration of key modules, including tasks, evaluation protocols (pointwise or pairwise), adversarial attacks (via Attacker Factory), defense techniques (via Defense Guards), judge prompt templates, and judge models.
RobustJudge aims to evaluate the robustness of LLM-as-a-Judge systems and generates quantitative metrics based on their judgment performance. As illustrated in Figure 2, RobustJudge follows a modular workflow that begins with selecting a benign input query from a specific tasks (e.g., translation, summarization, code generation)(§4.2). The response to this input, generated by a target LLM, is then processed through two parallel paths:
  • Benign path: The response is directly formatted using a specified Judge Prompt Template4.6) and sent to the LLM judges4.7) for evaluation.
  • Adversarial path: The same response is first manipulated by the Attacker Factory4.4), generating an adversarial variant. This response can optionally pass through the Defense Guard4.5) for mitigation, before being evaluated by the same LLM judge.
The LLM-as-a-Judge then evaluates both benign and adversarial responses, assigning scores (for pointwise evaluation protocol) or pairwise preferences (for pairwise evaluation protocol)(§4.3). In parallel, the responses are analyzed by a Content Quality Evaluator4.8), which provides content quality scores using automatic metrics, such as BLEURT for text tasks or CodeBLEU for code-related tasks. Finally, the results from both paths are consolidated in the Metrics module (§4.9), where RobustJudge quantifies robustness based on the judgments on the benign and adversarial responses and their content quality scores.
4.2 Tasks
In this paper, we evaluate LLM-as-a-Judge on 8 tasks. These tasks fall into 3 categories, e.g., text-focused, code-oriented and knowledge-intensive evaluation tasks. The details of these tasks are summarized in Table I.
Text-focused Tasks. This category aims to evaluate the robustness of judgment on fundamental text processing tasks. Specifically, we include two tasks, e.g., machine translation and text summarization.
Machine Translation (T1). This task focuses on translating text between languages while maintaining meaning, fluency, and accuracy. We adopt 6 language pairs from FLORES-200 benchmark [41], including high-resource pairs such as Chinese-English, German-English, as well as the low-resource Yoruba-English, all evaluated in both translation directions.
Text Summarization (T2). Text Summarization distills lengthy documents into concise, factual summaries. We draw on English news articles from the CNN/DailyMail dataset [42]. Each sample in this dataset contains an original article and its reference summary.
Code-oriented Tasks. Processing and generating code are essential capabilities for assessing the performance of both LLMs and LLM-as-a-Judge systems. To comprehensively evaluate judgment in the code domain, we include three representative tasks: code translation, code summarization, and code generation.
Code Translation (T3). Code Translation involves converting source code from one programming language to another while preserving its original functionality. For this evaluation, we use the Code-TransOcean dataset [42], focusing on the subset with C#-Java language pairs.
Code Summarization (T4). This task involves generating concise natural language descriptions that capture the functionality of given code snippets. For our evaluation, we utilize the Python code summarization subset from the CodeXGLUE benchmark  [43].
Code Generation (T5). The code generation task focuses on producing executable code based on a natural language specification or a function signature. Our evaluation is based on the samples from the CodeSearchNet dataset [44], specifically targeting Python function generation.
TABLE I: Judge tasks used in our dataset.
ID Category Task Name #Nums. Source Dataset
T1 Text Text Translation 30 Flores-200[41]
T2 Text Summarization 20 CNN/DailyMail[42]
T3 Code Code Translation 30 CodeTransOcean[45]
T4 Code Generation 20 CodeSearchNet[44]
T5 Code Summarization 20 CodeXGLUE[43]
T6 Knowledge Mathematics 20 LiveBench[46]
T7 Logical Reasoning 20 LiveBench[46]
T8 Knowledge Recall 28 MMLU-Pro[47]
Knowledge-intensive Tasks. We extend the evaluation of the LLM-as-a-Judge system’s robustness in knowledge-intensive scenarios. Following the setup in JudgeBench [8], we select three core areas, i.e., logical reasoning, mathematical problem solving, and professional knowledge recall.
logical Reasoning (T6). We incorporate benchmarks from Big-Bench Hard [48] and classic puzzles (e.g., the Zebra Puzzle) to examine the judgment robustness on abstract, compositional, and logical inference skills. These tasks measure the model’s ability to perform multi-step reasoning and handle complex constraint-based problems.
Mathematical Problem Solving (T7). Our mathematical evaluation suite comprises problems from AMC12 and USAMO competitions, including topics such as algebra, geometry, and combinatorics. These problems assess the model’s ability to perform numerical reasoning and formal proof strategies.
Professional Knowledge Recall (T8). We employ MMLU-Pro [47], an enhancement of the original MMLU [49]. MMLU-Pro contains college-level, multiple-choice questions across 14 professional disciplines, e.g., Physics, Chemistry, and Law.
4.3 Evaluation Protocol
LLM-as-a-Judge systems typically follow one of two evaluation protocols: pointwise or pairwise. In this study, we consider both protocols to assess the robustness of LLM-based judges under adversarial conditions.
Pointwise Evaluation. This protocol evaluates the quality of a single response. The LLM judge is prompted to assign an integer score $s$ (typically ranging from 1 to 10) that reflects the overall quality of the response within the given context. Formally, the evaluation process is represented as:
$s=M(P_{1},r)$ (1)
where $M$ denotes the judge model, $P_{1}$ denotes the pointwise prompt template and $r$ is the response being evaluated.
Pairwise Evaluation. This protocol compares two candidate responses and determines which one is preferred in the given context. The model outputs a preference for either response $r_{a}$ or $r_{b}$, depending on their relative quality. Formally, this can be expressed as:
$p=M(P_{2},r_{a},r_{b})$ (2)
where $P_{2}$ is the pairwise prompt template, and $r_{a}$ and $r_{b}$ are the two responses under evaluation, $p\in\{r_{a},r_{b}\}$. Prior research on LLM-as-a-Judge systems [5] has shown that the input order of the candidate responses can affect the evaluation outcome. To account for this bias, evaluations are typically conducted using both input orders, with the results defined as:
$p_{+}=M(P_{2},r_{a},r_{b}),\quad p_{-}=M(P_{2},r_{b},r_{a})$ (3)
4.4 Attacker Factory
Attacker factory is responsible for generating adversarial responses $r_{m}$ by applying various attack techniques designed to probe the vulnerabilities of LLM-as-a-Judge systems. These attacks are broadly categorized into heuristic-based and optimization-based methods, depending on how adversarial attacks are constructed. A summary of heuristic-based attack methods is listed in Table VIII (appendix), and optimization-based attacks can be found in Table IX (appendix).
Heuristic-based Attacks. These attacks are based on manually crafted prompts or linguistic manipulations that exploit LLM weaknesses in instruction following, context interpretation, or reasoning flow. Unlike optimization-based approaches, these attacks aim to alter the judgment outcome through direct prompt manipulations. Formally, a heuristic-based attack is defined as:
$r_{m}=A(r)$ (4)
where $r$ denotes the original benign response, and $A$ is the transformation applied by the attack. Our evaluation includes the following 8 heuristic-based attacks: Naive Attack (H1) [30], Escape Characters (H2) [31], Context Ignoring (H3) [32, 39], Fake Completion (H4) [40], Fake Reasoning (H5) [12], Combined Attack (H6) [14], Empty (H7) and Long-Suffix (H8).
Optimization-Based Attacks. These attack methods employ automated search or optimization algorithms to progressively manipulate response $r$. They iteratively modify candidate responses based on judge scores or internal signals. This process is formalized as:
$r_{m}^{(i+1)}=A\left(r_{m}^{(i)}\right),\quad\text{for }i=0,1,\ldots,n$ (5)
where $r_{m}^{(0)}=r$ is the initial benign response, and $n$ is the number of iterations. The 7 optimization-based attacks in our evaluation include: AdvEval (O1) [10], PAIR (O2) [15], TAP (O3) [16], Cheating (O4) [11], GCG (O5) [33], AutoDAN (O6) [34], and Greedy (O7) [9].
4.5 Defense Guard
To evaluate the impact of defense techniques against adversarial attacks for LLM-as-a-Judge systems, we incorporate a Defense Guard module in RobustJudge. Inspired by PromptBench [14], this module includes both prevention-based and detection-based strategies, which are lightweight and applicable without modifying the LLM judges. Further technical details are provided in Appendix  IX.
Prevention-based Defenses. These methods aim to enhance the actual input $I$ to the judge models by modifying the prompt $P$ and/or the response $r$ to prevent attacks from taking effect. This can be expressed as:
$I=D(P,r)$ (6)
where $D$ denotes the transformation function, such as re-tokenizing inputs or reformatting the prompt. We evaluate the following prevention methods: Retokenization (D1) [17], Delimiters (D2) [40], Sandwich Prevention (D3) [35], Instructional Prevention (D4) [50].
Detection-based Defenses. These methods aim to identify and filter out adversarial inputs before they reach the judge model. If an input is flagged as adversarial based on a predefined metric $f(\cdot)$, the evaluation is skipped. Formally:
$I=(P,r),\quad\text{iff }f(P,r)>\tau$ (7)
where $f(\cdot)$ outputs a scalar confidence score and $\tau$ is a detection threshold. We consider three detection strategies: PPL (D5) [37], Windowed PPL (D6) [17], Naive LLM-based (D7) [18].
4.6 Judge Prompt Template
In the context of LLM-as-a-Judge, a variety of judge prompt templates have been developed to guide LLMs in evaluating model outputs. To explore how these templates influence the robustness of LLM-as-a-Judge systems, we begin by collecting 3 widely adopted judge prompt templates, e.g., Vanilla Prompt[51], Arena-Hard Prompt[52], and Google Vertex Prompt[53] from JudgeBench [8]. We evaluate their performance under various adversarial attacks to examine their contribution to the robustness of LLM-as-a-Judge systems.
We hypothesize that the choice of judge prompt template plays a critical role in shaping the judge system’s robustness. Therefore, identifying more effective templates is essential for improving their robustness. To explore this further, we extend our analysis to include 6 additional commonly used prompt templates: PandaLM, Prometheus 2, JudgeLM, Auto-J, Skywork, and ChatEval. All these 9 instances collectively form a representative set of judge prompt templates.
We observe that these templates share a common structural format composed of key functional components. Based on this observation, we conduct a component-level analysis by decomposing each prompt into a set of 6 essential components that define the judge’s behavior.
  • Role Specification (RS): Defines the role of the LLM, such as an impartial judge, expert evaluator, or critic.
  • Evaluation Instruction (EI): Instructs the LLM to explicitly generate the score for a response (pointwise) or compare two candidate responses (pairwise).
  • Independent Answer Generation (IAG): Requires the LLM to generate its own answer prior to evaluating the given outputs.
  • Evaluation Criteria (EC): Specifies the factors or metrics for judgment, e.g., helpfulness, relevance, accuracy, fluency, and creativity.
  • Explanation Requirement (ER): Requires the LLM to provide an explanation for its judgment.
  • Rating Format (RF): Defines the output format of the judgment.
TABLE II: Summary of key components of judge prompt template.
Prompt RS EI IAG EC ER RF
Vanilla
Arena-Hard
Google Vertex
PandaLM
Prometheus 2
JudgeLM
Auto-J
Skywork
ChatEval
The summary of the decomposed judge prompts is provided in Table II.
Building on the identified components, we aim to determine the optimal prompt template configuration using an optimization-based approach. Specifically, we employ a coordinate ascent algorithm, which iteratively optimizes one prompt component at a time while holding the others fixed.
This process continues until convergence, yielding a prompt configuration that maximizes evaluation accuracy and robustness. We present the coordinate ascent algorithm in Algorithm 1.
Algorithm 1 Coordinate Ascent for Judge Prompt Optimization
0:  Prompt components $\mathcal{C}=\{C_{1},C_{2},\dots,C_{n}\}$, initial configuration $\mathbf{c}^{(0)}$, evaluation function $\mathcal{E}(\cdot)$
0:  Optimized prompt configuration $\mathbf{c}^{*}$
1:  Initialize $\mathbf{c}^{(0)}$ (e.g., randomly or using a known baseline)
2:  Set $t\leftarrow 0$
3:  repeat
4:     for each component $C_{i}\in\mathcal{C}$ do
5:        Fix all components except $C_{i}$
6:        Search for the best $c_{i}^{(t+1)}$ that maximizes $\mathcal{E}(\mathbf{c})$
7:        Update: $\mathbf{c}^{(t+1)}\leftarrow\mathbf{c}^{(t)}$ with $C_{i}=c_{i}^{(t+1)}$
8:     end for
9:     $t\leftarrow t+1$
10:  until convergence or maximum iterations reached
11:  return $\mathbf{c}^{(t)}$ as $\mathbf{c}^{*}$
4.7 LLM Judges
We consider 12 LLM-as-a-Judge systems with diverse characteristics across multiple dimensions: backbone architecture, model size, fine-tuning status, reasoning capability, and open-source availability.
Specifically, the evaluated LLM judges include proprietary systems such as GPT-4o and PAI-Judge, as well as open-source models built on LLaMA, Mistral, and Qwen backbones. Among them, several models (e.g., JudgeLM, PandaLM, Auto-J, Prometheus 2) have been explicitly fine-tuned for judgment tasks, while others (e.g., LLama-3.1, Openchat-3.5, Qwen2.5) remain zero-shot evaluators. We also include Ds-R1, a reasoning-oriented judge system built on Ds-V3, to explore whether enhanced reasoning capabilities influence judgment robustness and behavior. The features of the LLM judges are provided in Table XI (appendix).
4.8 Content Quality Evaluator
Our evaluation metrics (detailed in §4.9) are designed to capture not only the judgment accuracy of the LLM-as-a-Judge systems but also the content quality drift of the evaluated responses. To quantify this drift, we introduce a content quality evaluator module $Q$, which assesses how closely a target response aligns with a reference response. Formally, the content quality score $S_{e}$ is defined as:
$s_{e}=Q(x,r_{\text{ref}})$ (8)
where $x$ denotes the target response (either $r$ or $r_{m}$), and $r_{\text{ref}}$ is the corresponding reference response. We design specialized evaluators for different task categories, i.e., text-focused, code-oriented evaluation tasks, to ensure appropriate quality assessment.
Evaluator for Text-focused Tasks $Q_{\text{text}}$. For text-focused evaluation tasks such as summarization and translation, we assess the fluency, coherence, and relevance of the response using BLEURT [54], a state-of-the-art BERT-based model fine-tuned on human-labeled data. BLEURT computes semantic similarity between two responses, returning a score in the range (0,1). The reference responses $r_{\text{ref}}$ are drawn from benchmark datasets corresponding to each task.
Evaluator for Code-oriented Tasks $Q_{\text{code}}$. For code-oriented tasks, we introduce an evaluator to assess functional correctness, syntactic validity, and semantic accuracy. We adopt CodeBLEU  [55], a metric designed for programming languages. CodeBLEU extends the traditional BLEU score by incorporating code-specific properties, such as syntax structure, data-flow consistency, and token-level similarity. The final CodeBLEU score also lies in the range (0,1).
Evaluator for Knowledge-intensive Tasks $Q_{\text{know}}$. For knowledge-intensive tasks, such as logical reasoning and mathematical problem solving, semantic similarity measures fail to accurately reflect content quality. In these cases, the correspondence between the model response and the reference often involves abstract reasoning rather than surface-level similarity. As such, we disable the content quality evaluator for these tasks and assign a fixed score of $S_{e}=0$.
4.9 Evaluation Metrics
We define a set of metrics to quantify the impact of adversarial attacks and defense techniques.
  • Score Difference Rate (SDR): This metric applies to the pointwise evaluation protocol and measures the change in the output score assigned by the judge model before and after an adversarial attack (or defense). Formally, SDR is defined as:
    $\text{SDR}=\frac{1}{N}\sum_{i=1}^{n}(s_{t}^{(i)}-\hat{s}_{t}^{(i)})\times 0.1$ (9)
    where $N$ denotes the size of the test set. $\hat{s}_{t}^{(i)}$ is the original score for instance $i$, and $s_{t}^{(i)}$ is the score after attack or defense. The factor of 0.1 serves to normalize the $\hat{s}_{t}^{(i)}$ and $s_{t}^{(i)}$ value to the range (0, 1).
  • Improved Score Difference Rate (iSDR): Since adversarial modifications or defenses may enhance the quality of responses (leading to unintended score improvements) as a side effect, we propose iSDR to isolate the impact of manipulation from genuine content improvement. iSDR compensates for content quality changes by subtracting the corresponding content quality score change. Formally:
    iSDR $=\frac{1}{N}\sum_{i=1}^{n}\text{iSDR}^{(i)}$ (10)
    $=\frac{1}{N}\sum_{i=1}^{n}\left((s_{t}^{(i)}-\hat{s}_{t}^{(i)})\times 0.1-(s_{e}^{(i)}-\hat{s}_{e}^{(i)})\right)$
    where $\hat{s}_{e}^{(i)}$ and $s_{e}^{(i)}$ denote the content quality scores before and after the attack (or defense), respectively.
  • Attack Success Rate (ASR): ASR measures the proportion of successful adversarial attacks, defined differently for pointwise and pairwise evaluation protocols. Pointwise ASR (ASR) is defined as the proportion of test cases with positive iSDR, representing a manipulated judgment:
    $\text{ASR}=\frac{1}{N}\sum_{i=1}^{n}\mathbbm{1}(\text{iSDR}^{(i)}>0)$ (11)
    Pairwise ASR (P-ASR) considers both of the candidate orders and measures how often the judge fails to select the reference response $r_{\text{ref}}$ as the better choice.
[…]
    where $\mathbbm{1}(\cdot)$ is the indicator function, which returns 1 if the condition is true and 0 otherwise.
  • Accuracy, Recall, Precision, and F1: These metrics are used to evaluate the performance of detection-based defense techniques, which aim to distinguish between adversarial and benign responses. Specifically, the defense guard attempts to correctly identify adversarial responses (TP), preserve benign responses (TN), and avoid misclassifying a benign response as adversarial (FP) or failing to detect an adversarial response (FN). Based on these quantities, we compute Accuracy (ACC), Recall, Precision (Pre), and F1 Score.
TABLE III: Evaluation of adversarial attack for the responses generated by openchat-3.5 on translation tasks. Higher values indicate superior attack effectiveness for all metrics.
Judge Model Metric Attack Methods Average
H1 H2 H3 H4 H5 H6 H7 H8 O1 O2 O3 O4 O5 O6 O7
Openchat-3.5 ASR(%) 76.67 86.67 80.00 40.00 90.00 100.00 86.67 30.00 86.67 83.33 73.33 30.00 10.00 10.00 6.670 59.72
SDR 0.343 0.386 0.385 0.221 0.442 0.551 0.527 0.521 0.432 0.682 0.140 0.194 -0.008 -0.052 0.015 0.385
iSDR 0.082 0.125 0.124 -0.039 0.181 0.290 0.266 0.260 0.171 0.421 -0.121 -0.068 -0.269 -0.313 -0.246 0.061
P-ASR(%) 5.000 13.33 23.33 10.00 40.00 28.33 6.670 30.00 28.33 21.67 21.67 16.67 0.000 16.67 3.330 17.67
Qwen-2.5 ASR(%) 86.67 90.00 86.67 36.67 90.00 100.00 100.00 86.67 83.33 100.00 43.33 80.00 40.00 23.33 40.00 72.44
SDR 0.206 0.170 0.342 0.018 0.356 0.556 0.447 0.591 0.550 0.697 0.086 0.305 0.018 -0.060 0.051 0.289
iSDR 0.126 0.090 0.262 -0.061 0.276 0.476 0.367 0.464 0.470 0.617 0.006 0.225 -0.063 -0.140 -0.029 0.206
P-ASR(%) 6.670 11.67 43.33 21.67 28.33 60.00 3.330 1.670 33.33 25.00 15.00 3.330 3.330 1.670 8.330 17.78
Mistral-7B ASR(%) 70.00 83.33 56.67 76.67 89.66 100.00 93.33 76.67 90.00 86.67 93.33 42.86 40.00 23.33 40.00 70.83
SDR 0.110 0.134 0.089 0.323 0.446 0.565 0.633 0.359 0.575 0.676 0.580 0.160 -0.062 0.085 0.036 0.315
iSDR 0.035 0.059 0.014 0.248 0.374 0.490 0.558 0.284 0.473 0.602 0.506 0.090 -0.136 0.010 -0.039 0.238
P-ASR(%) 15.00 20.00 8.330 66.67 48.33 63.33 20.00 60.00 60.00 33.33 50.00 26.67 13.33 6.670 35.00 35.11
Llama-3.1-8B ASR(%) 63.33 60.00 56.67 60.00 60.00 90.00 86.67 80.00 70.00 90.00 56.67 56.67 43.33 36.67 56.67 65.11
SDR 0.069 0.010 -0.016 0.051 0.146 0.599 0.587 0.486 0.362 0.747 0.185 0.206 0.125 0.113 0.042 0.248
iSDR 0.072 0.013 -0.013 0.054 0.149 0.601 0.650 0.328 0.394 0.750 0.027 0.048 -0.033 -0.045 0.080 0.205
P-ASR(%) 10.00 20.00 51.67 46.67 46.67 71.67 5.000 10.00 20.00 51.67 45.00 3.330 10.00 40.00 23.33 30.33
Llama-3.3-70B ASR(%) 33.33 90.00 86.67 36.67 90.00 96.67 96.67 100.00 93.33 66.67 23.33 16.67 16.67 16.67 26.67 60.89
SDR 0.136 0.232 0.403 -0.008 0.282 0.516 0.532 0.678 0.432 0.443 -0.049 0.194 -0.140 -0.138 -0.010 0.234
iSDR -0.122 0.096 0.267 -0.144 0.146 0.380 0.395 0.542 0.513 0.307 -0.185 -0.208 -0.276 -0.274 -0.146 0.086
P-ASR(%) 1.670 3.330 28.33 3.330 8.330 3.330 0.000 0.000 11.67 0.000 0.000 0.000 0.000 0.000 3.330 4.220
5 Experiments
In this section, we leverage RobustJudge to systematically compare the effectiveness of different adversarial attack methods against the LLMs listed in Table XI [^1] We accessed all models via official APIs or hosted endpoints provided by Together.ai, HuggingFace, and OpenAI.. Our evaluation aims to answer the following research questions:
  • RQ1: What impact do different adversarial attacks and defense methods have on the LLM-based judges?
  • RQ2: How do the prompt templates and model choices affect the robustness of LLM-based judges?
  • RQ3: What vulnerabilities persist in black-box real-world deployments of LLM-based judges, as revealed by our empirical evaluations?
5.1 Attack and Defense Effectiveness (RQ1)
5.1.1 Attack Performance
We conduct all main experiments on Openchat-3.5 and present results on other judge models (e.g., Qwen-2.5) in Table XIII (appendix).
Impact of Advarsarial Attacks. We evaluate the effectiveness of various adversarial attack methods, and report their ASR, SDR, iSDR and P-ASR in Table III.
Heuristic-based attacks generally exhibit strong and stable effectiveness across all evaluated models. For instance, most heuristic attacks (H1–H7) consistently achieve high ASRs, frequently exceeding 80%. Notably, the Combined Attack (H6) achieves 100% ASR on Openchat-3.5, Qwen-2.5, and Mistral-7B, and above 90% on Llama-3.1-8B and Llama-3.3-70B. This trend is also reflected in SDR and iSDR, where H6 yields the highest SDR values, e.g., 0.5510 (Openchat-3.5), 0.5560 (Qwen-2.5), 0.5652 (Mistral-7B), 0.5985 (Llama-3.1-8B), and 0.5160 (Llama-3.3-70B). Additionally, H6 exhibits high impact across different underlying judge model. We attribute the effectiveness of H6 to its multi-faceted attack strategy, which integrates multiple manipulations, such as escape characters, context manipulation, and adversarial completions, into a single composite prompt. The attack succeeds if any one of its constituent mechanisms is effective, allowing its effects to accumulate. These results suggest that combining compatible attack techniques can lead to amplified adversarial effectiveness, highlighting the importance of defense mechanisms that address multi-facet threat scenarios.
Optimization-based attacks are also effective against LLM-as-a-Judge systems, particularly methods O1 (AdvEval), O2 (PAIR), and O3 (TAP). For instance, O1 achieves 86.67% ASR on Qwen-2.5, 90% ASR on Mistral-7B, and 93.33% on Llama-3.3-70B. Similarly, O2 attains 100% ASR on Qwen-2.5 and 90% on Llama-3.1-8B, while O3 also yields consistently high ASR scores across multiple models. In addition to ASR, we observe a similar trend on the other metrics such as SDR and iSDR, which further confirms the effects of these attacks. In contrast, other optimization-based approaches (O5–O7, e.g., GCG, AutoDAN, Greedy) result in lower ASRs (typically below 50%) and exhibit limited or even negative SDR/iSDR improvement. These results demonstrate their relatively weak optimization effectiveness and poor cross-model transferability. These results also reflect our experimental constraints. Given the high computational cost of optimization-based attacks, we restrict the number of iterations for each run. Although this may affect the final performance, it ensures practical feasibility and fair comparison.
$\bullet$ Finding 1: Existing LLM-as-a-Judge systems are still highly vulnerable to adversarial attacks, particularly methods Fake Reasoning (H5), Combined Attack (H6), Empty (H7) and AdvEval (O1), PAIR (O2), TAP (O3).
TABLE IV: Evaluation results across multiple tasks (T1–T8) and models under three representative attack methods: Combined Attack (H6), Fake Reasoning (H5), and Fake Completion (H4).
Type Task Judge Model Attack Methods
Combined Attack (H6) Fake Reasoning (H5) Fake Completion (H4)
ASR (%) SDR iSDR P-ASR (%) ASR (%) SDR iSDR P-ASR (%) ASR (%) SDR iSDR P-ASR (%)
Text T1 openchat-3.5 100.00 0.551 0.290 28.33 90.00 0.442 0.181 40.00 40.00 0.221 -0.040 10.00
Qwen2.5-7B 100.00 0.556 0.476 60.00 90.00 0.356 0.276 28.33 36.67 0.019 -0.061 21.67
Mistral-7B 100.00 0.565 0.491 63.33 89.66 0.446 0.372 48.33 76.67 0.323 0.248 66.67
LLama-3.1-8B 90.00 0.599 0.601 71.67 60.00 0.146 0.149 46.67 60.00 0.051 0.054 46.67
T2 openchat-3.5 90.00 0.491 0.258 45.00 85.00 0.438 0.204 90.00 85.00 0.466 0.232 67.50
Qwen2.5-7B 95.00 0.541 0.346 55.00 95.00 0.480 0.285 87.50 100.00 0.457 0.262 52.50
Mistral-7B 90.00 0.491 0.207 97.50 90.00 0.509 0.224 90.00 95.00 0.519 0.234 100.00
LLama-3.1-8B 80.00 0.370 0.150 75.00 80.00 0.365 0.145 65.00 90.00 0.373 0.154 35.00
Code T3 openchat-3.5 83.33 0.593 0.100 35.00 63.33 0.508 0.016 31.67 73.33 0.625 0.133 41.67
Qwen2.5-7B 93.33 0.648 0.288 0.00 70.00 0.438 0.078 56.67 56.67 0.411 0.051 61.67
Mistral-7B 83.33 0.734 0.163 98.33 83.33 0.740 0.169 80.00 83.33 0.700 0.129 96.67
LLama-3.1-8B 86.67 0.641 0.136 55.00 66.67 0.507 0.001 33.33 63.33 0.531 0.025 51.67
T4 openchat-3.5 85.00 0.716 0.118 10.00 65.00 0.698 0.100 50.00 80.00 0.723 0.125 60.00
Qwen2.5-7B 95.00 0.742 0.143 22.50 70.00 0.659 0.061 95.00 80.00 0.703 0.105 97.50
Mistral-7B 100.00 0.742 0.141 100.0 100.00 0.768 0.167 87.50 100.00 0.761 0.160 90.00
LLama-3.1-8B 95.00 0.742 0.160 87.50 80.00 0.653 0.072 60.00 75.00 0.653 0.071 77.50
T5 openchat-3.5 100.00 0.593 0.199 45.00 90.00 0.552 0.158 92.50 95.00 0.608 0.214 70.00
Qwen2.5-7B 35.00 0.602 0.008 57.50 45.00 0.613 0.018 57.50 45.00 0.581 -0.013 62.50
Mistral-7B 100.00 0.581 0.080 100.0 95.00 0.551 0.051 97.50 95.00 0.618 0.117 97.50
LLama-3.1-8B 100.00 0.581 0.264 82.50 95.00 0.485 0.168 92.50 95.00 0.553 0.236 60.00
Knowledge T6 openchat-3.5 70.00 0.255 0.229 29.46 75.00 0.242 0.216 27.68 85.00 0.264 0.238 35.00
Qwen2.5-7B 95.00 0.546 0.417 1.790 95.00 0.320 0.192 14.29 50.00 0.134 0.005 20.00
Mistral-7B 90.00 0.950 0.220 77.50 95.00 0.938 0.208 57.50 90.00 0.865 0.135 60.00
LLama-3.1-8B 100.00 0.998 0.306 35.00 75.00 0.824 0.133 7.500 65.00 0.772 0.081 17.50
T7 openchat-3.5 65.00 0.359 0.348 23.47 60.00 0.356 0.345 20.41 40.00 0.138 0.126 17.50
Qwen2.5-7B 95.00 0.369 0.342 1.020 45.00 0.170 0.143 32.65 60.00 0.097 0.069 10.00
Mistral-7B 85.00 0.900 0.223 85.00 70.00 0.768 0.091 55.00 70.00 0.760 0.083 57.50
LLama-3.1-8B 85.00 0.850 0.163 62.50 65.00 0.765 0.078 10.00 60.00 0.756 0.069 30.00
T8 openchat-3.5 85.71 0.445 0.437 17.86 89.21 0.413 0.406 22.40 50.00 0.120 0.112 23.21
Qwen2.5-7B 35.71 0.183 0.178 13.33 10.71 0.048 0.043 38.33 $0$0.004 -0.008 35.71
Mistral-7B 53.57 0.718 0.252 73.21 50.00 0.650 0.184 44.64 39.29 0.455 -0.011 64.29
LLama-3.1-8B 85.71 0.921 0.301 41.07 67.86 0.853 0.233 21.43 71.43 0.856 0.236 41.07
Attacks on different Tasks Categories. We conduct evaluation on a variety of tasks, e.g., text-focused, code-oriented and knowledge-intensive tasks, to examine how vulnerabilities to adversarial attacks varies by task categories. The results are presented in Table IV. Additional results on other tasks are provided in Tables XII, XIV, XV, XVI, XVII, XVIII, and XIX in the appendix.
We observe substantial differences in ASR across these tasks. Text-focused tasks (T1–T2), such as machine translation and summarization, exhibit the highest vulnerability. For example, under the Combined Attack, all judge models demonstrate ASRs exceeding 90% on machine translation (T1). In contrast, knowledge-intensive tasks, such as mathematical problem solving (T7) and knowledge recall (T8) show significantly greater robustness. For instance, Qwen2.5-7B achieves an ASR of only 0% on T8 under the Fake Completion attack, and just 45% ASR on T7 under Fake Reasoning attack (H5). This observation highlights the need for task-aware defense mechanisms, especially when deploying LLM-as-a-Judge systems on text-focused tasks.
We believe these discrepancies stem from fundamental differences in task complexity and evaluation depth. Text-focused tasks often rely on surface-level features, such as fluency or coherence, which adversarial attacks can easily manipulate. As a result, LLM judges may be misled by outputs that appear syntactically sound but are semantically or factually incorrect. In contrast, knowledge-intensive tasks require deeper semantic understanding, logical reasoning, or domain-specific knowledge accuracy. These tasks are less vulnerable to superficial prompt modification. This analysis highlights the importance of task-aware defense strategies, particularly when deploying LLM-as-a-Judge systems on text-focused tasks.
$\bullet$ Finding 2: Tasks that emphasize surface-level features (e.g., text-forcused task) exhibit greater vulnerability than knowledge-intensive tasks, which require deeper semantic understanding and reasoning.
Comparison of Judge Protocals. In this section, we compare two evaluation protocols used on LLM-as-a-Judge systems: pointwise scoring and pairwise comparison. The corresponding results are shown in Table III and Figure 5.
Our findings show that while many adversarial strategies achieve high ASR under pointwise scoring, their P-ASR drops considerably in the pairwise comparison setting. For instance, Combined Attack (H6) achieves a perfect ASR of 100% against Openchat-3.5, yet its P-ASR is only 28.33%. Similar trends are observed for Qwen2.5-7B-Instruct (ASR: 100%, P-ASR: 60.00%) and Mistral-7B (ASR: 100%, P-ASR: 63.33%). This discrepancy can be attributed to the nature of the pairwise protocol, which provides the judge model with an additional reference response for comparison. This allows the model to contrast the manipulated response with a stable reference, thereby reducing the attack’s impact.
Despite the absolute differences in ASR and P-ASR, the relative effectiveness of attacks remains consistent for both protocols. Combined Attack (H6) and Fake Reasoning (H5) consistently rank among the most effective attacks in both ASR and P-ASR against all judge models, as shown in Table III. Conversely, optimization-based attack GCG (O5) and Auto-DAN (O6) exhibit limited impact on Openchat-3.5 (e.g., 0% and 6.67% ASR, respectively) under pointwise protocol, and correspondingly low P-ASRs (e.g., P-ASR <10% in most cases) under the pairwise protocol.
$\bullet$ Finding 3: The relative effectiveness of attacks remains consistent across pointwise and pairwise protocols, i.e., attacks that achieve lower (or higher) ASR in one protocol also tend to yield lower (or higher) ASR in another.
TABLE V: Evaluation of prevention-based defense methods (D1-D4) against adversarial attacks.
Defense Metric Attack Methods Average
H1 H2 H3 H4 H5 H6 H7 H8 O1 O2 O3 O4 O5 O6 O7
D1 ASR(%) 0.000 0.000 0.000 0.000 3.330 100.00 90.00 3.330 3.330 3.330 3.330 23.33 6.670 6.670 6.670 16.67
SDR -0.055 -0.084 -0.053 -0.126 -0.018 0.845 0.629 0.171 -0.041 -0.023 -0.031 0.054 -0.100 -0.036 -0.021 0.074
iSDR -0.642 -0.698 -0.646 -0.574 -0.633 0.281 0.465 -0.336 -0.508 -0.417 -0.386 -0.150 -0.441 -0.405 -0.354 -0.363
D2 ASR(%) 73.33 80.00 80.00 53.33 90.00 96.67 93.33 40.00 76.67 76.67 80.00 33.33 6.670 10.00 20.00 60.67
SDR -0.070 -0.020 -0.016 -0.317 0.097 0.117 -0.108 -0.289 -0.049 -0.082 -0.126 -0.594 -0.634 -0.634 -0.517 -0.216
iSDR 0.066 0.088 0.112 -0.043 0.204 0.275 0.450 -0.074 0.207 0.245 0.241 -0.075 -0.253 -0.281 -0.128 0.069
D3 ASR(%) 76.67 86.67 86.67 46.67 96.67 100.0 86.67 50.00 83.33 83.33 90.00 33.33 0.000 6.670 23.33 63.33
SDR -0.023 0.012 0.048 -0.329 0.106 0.147 -0.064 -0.215 0.001 -0.058 -0.040 -0.594 -0.768 -0.655 -0.533 -0.198
iSDR 0.102 0.110 0.166 -0.065 0.203 0.295 0.485 -0.010 0.246 0.260 0.318 -0.085 -0.397 -0.312 -0.154 0.077
D4 ASR(%) 80.00 80.00 80.00 46.67 93.33 100.0 90.00 43.33 80.00 76.67 73.33 26.67 13.33 10.00 26.67 61.33
SDR -0.035 -0.033 -0.043 -0.329 0.045 0.119 -0.124 -0.287 -0.051 -0.126 -0.161 -0.662 -0.690 -0.713 -0.588 -0.245
iSDR 0.114 0.089 0.099 -0.042 0.166 0.291 0.449 -0.058 0.218 0.216 0.220 -0.130 -0.296 -0.347 -0.185 0.054
TABLE VI: Evaluation of detection-based defense methods (D5–D7) against adversarial attacks.
Defense Metric Attack Methods Average
H1 H2 H3 H4 H5 H6 H7 H8 O1 O2 O3 O4 O5 O6 O7
D5 ACC(%) 55.00 53.33 51.67 50.00 48.33 50.00 71.67 23.64 46.67 31.67 28.33 51.67 51.67 55.00 56.67 48.35
Recall(%) 66.67 63.33 60.00 56.67 53.33 56.67 100.00 0.00 50.00 20.00 13.33 60.00 60.00 66.67 70.00 53.11
Pre(%) 54.05 52.78 51.43 50.00 48.48 50.00 63.83 0.00 46.88 26.09 19.05 51.43 51.43 54.05 55.26 44.98
F1 0.597 0.576 0.554 0.531 0.508 0.531 0.779 0.000 0.484 0.226 0.157 0.554 0.554 0.597 0.618 0.484
D6 ACC(%) 53.33 55.00 55.00 50.00 46.67 48.33 70.00 20.00 46.67 45.00 50.00 50.00 70.00 53.33 55.00 51.22
Recall(%) 66.67 70.00 70.00 60.00 53.33 56.67 100.00 0.00 53.33 50.00 60.00 60.00 100.00 66.67 70.00 62.44
Pre(%) 52.63 53.85 53.85 50.00 47.06 48.57 62.50 0.00 47.06 45.45 50.00 50.00 62.50 52.63 53.85 48.66
F1 0.588 0.609 0.609 0.545 0.500 0.523 0.769 0.000 0.500 0.476 0.545 0.545 0.769 0.588 0.609 0.545
D7 ACC(%) 61.67 60.00 70.00 71.67 50.00 61.67 75.00 58.33 40.00 58.33 63.33 81.67 83.33 73.33 76.67 65.67
Recall(%) 56.67 53.33 73.33 76.67 33.33 56.67 83.33 50.00 13.33 50.00 60.00 96.67 100.00 80.00 86.67 64.71
Pre(%) 62.96 61.54 68.75 69.70 50.00 62.96 71.43 60.00 28.57 60.00 64.29 74.36 75.00 70.59 72.22 65.49
F1 0.597 0.571 0.710 0.730 0.400 0.596 0.769 0.545 0.182 0.545 0.621 0.841 0.857 0.750 0.788 0.627
5.1.2 Defense Performance
We evaluate 7 representative defenses methods from both prevention-based and detection-based paradigms. These include: retokenization[17] (D1), delimiter insertion (D2), sandwich prompting (D3), instruction augmentation (D4), perplexity filtering (PPL) (D5), windowed PPL (WinPPL) (D6), and a naive LLM-based detector (D7). Since prevention-based and detection-based methods are based on different mechanism and evaluated on different metrics, we report their results separately in Table V and Table VI, respectively.
Prevention-based Defenses (D1–D4). Overall, prevention-based methods demonstrate moderate effectiveness across most attacks. However, they remain largely ineffective against stronger adversarial attacks, such as Combined Attack (H6) and AdvEval (H7). Among these methods, Retokenization (D1) yields the best average defense performance, achieving a reduced ASR of 16.67% and an iSDR of -0.394, significantly outperforming the other methods (D2–D4), which exhibit ASRs exceeding 60% and positive iSDR values above 0.03. Despite this relative success, prior findings [17] indicate that Retokenization (D1) can degrade model performance on benign inputs. This degradation arises from altered tokenization patterns, potentially affecting how the model interprets the content and leading to inaccurate evaluations. Thus, while Retokenization offers strong adversarial robustness, it introduces a trade-off between defense effectiveness and benign input performance.
Detection-based Defenses (D5–D7) Among detection-based methods, perplexity-based filters (PPL and WinPPL) generally underperform. They achieve relatively low accuracy (48.35% and 51.22%), recall (53.11% and 62.44%), precision (45.01% and 48.66%), and F1 scores (0.518 and 0.545), indicating limited capability in distinguishing adversarial inputs. In contrast, the naive LLM-based detector achieves the highest detection performance across all metrics, with an accuracy (65.67%), recall (64.71%), precision (65.49%) and F1 score (62.68%), outperforming all other prevention-based defenses, particularly against complex heuristic-based manipulations. Its effectiveness stems from leveraging a powerful LLM to identify subtle adversarial patterns in responses. However, this approach incurs significant computational overhead, as it requires querying the LLM detector for each input response.
$\bullet$ Finding 4: Retokenization and the Naive LLM-based Detector are two of the most effective defense methods against adversarial attacks. However, their effectiveness comes with trade-offs: Retokenization may reduce performance on benign inputs due to altered tokenization patterns, while the LLM-based Detector introduces substantial computational overhead.
5.2 Impact of Prompts and Model Choices (RQ2)
TABLE VII: Comparison of robustness across different judge prompt templates on Text Translation tasks(T1).
Higher ASR (weaker robustness)
Lower ASR (stronger robustness)
Judge Model H4 H5 H6 Avg
openchat-3.5 12.92% 10.00% 10.00% 10.97%
Qwen2.5-7B 5.000% 15.83% 14.16% 11.99%
Mistral-7B 37.50% 40.83% 42.50% 40.28%
LLama-3.1-8B 6.670% 25.00% 11.67% 13.00%
(a) Vanilla Prompt
Judge Model H4 H5 H6 Avg
openchat-3.5 3.330% 0.000% 2.910% 2.080%
Qwen2.5-7B 39.17% 0.830% 6.250% 15.42%
Mistral-7B 96.25% 1.670% 87.08% 61.67%
LLama-3.1-8B 0.000% 0.830% 2.500% 1.110%
(b) Arena-Hard Prompt
Judge Model H4 H5 H6 Avg
openchat-3.5 7.080% 1.670% 0.420% 3.060%
Qwen2.5-7B 27.50% 0.000% 1.670% 9.760%
Mistral-7B 49.17% 33.33% 50.42% 44.31%
LLama-3.1-8B 73.33% 0.000% 55.00% 42.78%
(c) Google Vertex Prompt
Judge Model H4 H5 H6 Avg
openchat-3.5 14.17% 0.830% 3.330% 6.110%
Qwen2.5-7B 2.500% 0.000% 8.330% 3.610%
Mistral-7B 14.17% 0.000% 38.33% 17.50%
LLama-3.1-8B 1.670% 1.670% 0.000% 1.110%
(d) Optimized Prompt (ours)
5.2.1 Impact of Judge Prompt Template
We examine the impact of prompt templates on the robustness of LLM-as-a-Judge systems. Specifically, we evaluate three widely used prompt templates, e.g., Vanilla Prompt, Arena-Hard Prompt and Google Vertex Promp, across four judge models on the translation task under three adversarial attacks (H4–H6). In this evaluation, a higher ASR indicate weaker robustness.
The results are shown in Table VII (a-c). We observe that LLM-as-a-Judge systems are highly sensitive to the choice of prompt template. The same judge model can exhibit significantly different ASR values depending on the template used. For example, under H4, Mistral-7B yields an ASR of 37.50% with the Vanilla prompt, which increases to 96.25% with the Arena-Hard prompt. Similarly, LLama-3.1-8B displays a wide range in ASR, from 6.67% ASR (Vanilla) to 73.33% (Google Vertex) under H4, further highlighting this sensitivity.
To improve robustness, we apply the coordinate ascent strategy described in Algorithm 1 to derive an optimized prompt. Initializing from the Arena-Hard template, we iteratively refine individual prompt components based on ASR under the H4 attack. The optimization is performed on the translation task, with chinese2english language pair. We conduct three iterations of coordinate ascent using Openchat-3.5 as the target model. The optimized prompt is then evaluated for generalization across five additional language pairs within the translation task, and the transferability under H5 and H6 attacks.
Compared with the baselines, our optimized prompt achieves superior robustness (Table VII (d)). On Openchat-3.5, it reduces the average ASR across H4–H6 to 6.11%, outperforming Arena-Hard (20.08%) and Google Vertex (23.06%), while also maintaining comparably low ASR on Qwen-2.5 and LLaMA-3.1-8B. Notably, for Mistral-7B, it brings the ASR under H4 down from 96.25% (Arena-Hard) to only 14.17%, demonstrating a substantial improvement in robustness. The optimized prompt template is presented in Table X (appendix).
Importantly, unlike prior work that relies on extensive prompt search or model fine-tuning, our method performs optimization in a highly cost-efficient manner, targeting a single attack (H4), a single task (translation), and one language pair (chinese2english). Despite this constrained budget, the optimized prompt generalizes well on additional attacks (H5, H6) and models (Openchat, Qwen, Mistral, LLaMA), highlighting the strong generalizability and transferability of our approach. This exhibits that lightweight, targeted prompt modifications can yield broad defensive benefits and robustness for LLM-as-a-Judge systems.
$\bullet$ Finding 5: The robustness of LLM-as-a-Judge systems is highly sensitive to the choice of prompt template.
$\bullet$ Finding 6: We propose a low-cost yet highly transferable prompt template optimization method that improves robustness. The optimized template consistently outperforms existing prompt template.
(a) Empty Attack
(b) Escape Attack
(c) Ignore Attack
(d) Combined Attack
(e) Completion Attack
(f) Naive Attack
(g) Reasoning Attack
(h) Average
Figure 3: Evaluate the robustness of different judge models on 7 attacks.
5.2.2 Impact of Judge Model
We evaluate a range of judge models, including GPT-4o, Openchat, JudgeLM-7/13B, Prometheus-7B, DeepSeek-R1, PandaLM-7B, and AutoJ-13B, comparing their robustness against various adversarial attacks. The results are summarized in Figure  3. GPT-4o, considered to be the defacto LLM-as-a-Judge system, demonstrates strong robustness across different adversarial attacks, achieving an ASR of 71.26% and an iSDR of 0.1235 (Figure 3h). These results validate the current practice of using GPT-4o as a judge model and suggest that it offers a relatively reliable evaluation setup. Notably, JudgeLM-13B, a fine-tuned model specifically trained for judgment tasks, achieves the lowest ASR (69.00%) among all evaluated models and exhibits consistently strong robustness across different attack strategies. This highlights the effectiveness of task-specific fine-tuning in improving the robustness of LLM-as-a-Judge systems. Given its open-source availability and lower computational overhead compared to GPT-4o, JudgeLM-13B presents a promising and cost-effective alternative for reliable LLM-as-a-Judge system. In contrast, reasoning-oriented models such as DeepSeek-R1 have been hypothesized to offer superior robustness in judgment tasks due to their enhanced reasoning capabilities. However, our results show that DeepSeek-R1 achieves only moderate robustness, with an ASR of 75.20% and an iSDR of 0.5643, indicating no significant advantage over other models in adversarial scenarios.
$\bullet$ Finding 7: Judge-tuned models demonstrate the strongest robustness against adversarial attacks due to alignment-focused fine-tuning.
$\bullet$ Finding 8: Reasoning-focused models (e.g., DeepSeek-R1) are more susceptible to severe scoring errors when attacks succeed, indicating weaker reliability under adversarial conditions.
5.3 Real-World Case Study (RQ3)
We conduct experiments on a real-world industrial LLM-as-a-Judge platform, i.e., PAI-Judge [^2] https://pai.console.aliyun.com/#/ai-service/judge/overview. This platform provides an overall evaluation score for each input, along with a set of subscores reflecting multiple assessment dimensions, such as accuracy, fluency, and consistency. According to its official documentation, PAI-Judge also offers a premium version PAI-Judge-Plus, which is built on a larger model, and is reported to deliver superior judgment performance.
Our initial evaluation targets PAI-Judge using pointwise protocol and includes both heuristic-Based attacks (e.g., Naive Attack, Escape Characters Attack, Context Ignore Attack, Fake Completion Attack, Fake Reasoning, Combined Attack (H1-H8)), and optimization-based attacks (PAIR (O2)). To ensure compatibility with the platform, we carefully adapt the prompt template to match the PAI-Judge’s format. However, our results show that these conventional adversarial strategies had minimal effect on the platform’s evaluation outcome. This suggests that PAI-Judge incorporates effective internal defense mechanisms, making it relatively robust to standard adversarial manipulations. Sample evaluation results are presented in Table VIII (appendix).
(a) PAI-Judge
(b) PAI-Judge-Plus
Figure 4: Comparison of Avg Subscore and Overall Score for PAI-Judge and PAI-Judge-Plus.
To further probe the system’s robustness, we extended the PAIR-optimized adversarial prompts by appending long-suffix manipulations (lengthy irrelevant or distracting content). Surprisingly, this composite attack results in high scores on both PAI-Judge (a) and PAI-Judge-Plus (b), despite that these adversarial responses should have received near-zero scores (Figure 4). We further investigate the attack by gradually increasing the length of the appended suffix and observe a rising trend in effectiveness. While PAI-Judge-Plus demonstrates some resistance, its scoring eventually increases significantly once the suffix length approaches 2000 letters. These results reveal a critical loophole in PAI-Judge and PAI-Judge-Plus platform.
Using our RobustJudge framework, we are able to systematically evaluate and uncover this vulnerability in a real-world judge system. These findings demonstrate the practical utility of RobustJudge in identifying hidden flaws and guiding the development of more robust LLM-as-a-Judge systems. We responsibly disclosed this vulnerability to the PAI-Judge platform team prior to the release of our results.
$\bullet$ Finding 9: Conventional adversarial attacks were largely ineffective against the industrial PAI-Judge platform.
$\bullet$ Finding 10: We identify a loophole in both PAI-Judge and PAI-Judge-Plus, where combining PAIR-optimized adversarial inputs with long-suffix can compromise the platform’s judgment reliability.
6 Conclusion
This work presents the first scalable and fully automated framework to evaluate the robustness and reliability of LLM-as-a-Judge systems across multiple attack scenarios. We systematically benchmarked state-of-the-art LLM-based evaluators under various adversarial settings and found that they are vulnerable to manipulation, often producing biased or incorrect judgments when exposed to crafted inputs. Based on these findings, we conclude that current LLM-as-a-Judge systems are not yet sufficiently robust for secure deployment in critical evaluation tasks, and highlight the need for future research to strengthen their resilience against adversarial attacks. Our framework and benchmarks provide a valuable tool for the community to assess and track the progress of LLM-based evaluators in adversarial robustness.

### v2

@@
 Abstract
-Large Language Models (LLMs) have demonstrated remarkable intelligence across various tasks, which has inspired the development and widespread adoption of LLM-as-a-Judge systems for automated model testing, such as red teaming and benchmarking.
-However, these systems are susceptible to adversarial attacks that can manipulate evaluation outcomes, raising concerns about their robustness, hence trustworthiness.
-Existing evaluation methods adopted by LLM-based judges are often piecemeal and lack a unified framework for comprehensive assessment.
-Furthermore, the prompt template and model selections for improving judge robustness have been rarely explored, and their performance in real-world settings remains largely unverified.
+Large Language Models (LLMs) have demonstrated exceptional capabilities across diverse tasks, driving the development and widespread adoption of LLM-as-a-Judge systems for automated evaluation, including red teaming and benchmarking.
+However, these systems are susceptible to adversarial attacks that can manipulate evaluation outcomes, raising critical concerns about their robustness and trustworthiness.
+Existing evaluation methods for LLM-based judges are often fragmented and lack a unified framework for comprehensive robustness assessment.
+Furthermore, the impact of prompt template design and model selection on judge robustness has rarely been explored, and their performance in real-world deployments remains largely unverified.
 To address these gaps, we introduce RobustJudge, a fully automated and scalable framework designed to systematically evaluate the robustness of LLM-as-a-Judge systems.
-RobustJudge investigates the impact of attack methods and defense strategies (RQ1), explores the influence of prompt template and model selection (RQ2), and assesses the robustness of real-world LLM-as-a-Judge applications (RQ3).
-Our main findings are that 1) LLM-as-a-Judge systems are still vulnerable to a range of adversarial attacks, including Combined Attack and PAIR, while defense mechanisms such as Re-tokenization and LLM-based Detectors offer improved protection.
-2) Robustness is highly sensitive to the choice of prompt template and judge models.
-Our proposed prompt template optimization method can improve robustness, and JudgeLM-13B demonstrates strong performance as a robust open-source judge.
-3) Applying RobustJudge to Alibaba’s PAI platform reveals previously unreported vulnerabilities.
-The source code of RobustJudge is provided at https://github.com/S3IC-Lab/RobustJudge.
-1 Introduction
-Large Language Models (LLMs), such as OpenAI’s GPT-4o [1], Google’s Gemma2 [2], Meta’s Llama 3 [3], and QwenLM’s Qwen2.5 [4], have achieved remarkable proficiency across a wide range of tasks.
-Built on extensive training data and Transformer architectures, these models demonstrate advanced capabilities in general-purpose intelligence, including natural language understanding, text generation, and complex problem-solving.
+Specifically, RobustJudge investigates the effectiveness of 15 attack methods and 7 defense strategies across 12 models (RQ1), examines the impact of prompt template design and model selection (RQ2), and evaluates the security of real-world deployments (RQ3).
+Our study yields three key findings: (1) LLM-as-a-Judge systems are highly vulnerable to attacks such as PAIR and combined attacks, while defense mechanisms such as re-tokenization and LLM-based detectors can provide enhanced protection; (2) robustness varies substantially across prompt templates (up to 40%); (3) deploying RobustJudge on Alibaba’s PAI platform uncovers previously undiscovered vulnerabilities.
+These results offer practical insights for building trustworthy LLM-as-a-Judge systems.
+I Introduction
+Large Language Models (LLMs), such as OpenAI’s GPT-4o 2024gpt4 , Google’s Gemma2 gemma2 , Meta’s Llama 3 llama3 , and QwenLM’s Qwen2.5 2025qwen , have achieved remarkable proficiency across a wide range of tasks.
+Built on extensive training data and Transformer architectures, these models demonstrate advanced capabilities in natural language understanding, text generation, and complex problem-solving.
 To leverage these capabilities while minimizing human effort and mitigating human biases in LLM evaluation, the concept of LLM-as-a-Judge has been introduced.
-LLM-as-a-Judge [5] aims to automate the assessment of LLM-generated content, providing a scalable and objective alternative to human evaluation.
-This approach has gained widespread adoption, becoming a defacto evaluation method for assessing LLM performance across various domains, including software engineering[6], domain-specific knowledge assessment[7], and mathematical reasoning[8].
+LLM-as-a-Judge judgeFirst aims to automate the assessment of LLM-generated content, providing a scalable and objective alternative to human evaluation.
+This approach has gained widespread adoption, becoming an extensive evaluation method for assessing LLM performance across various domains, including software engineering codeJudge , domain-specific knowledge assessment legalJudge , and mathematical reasoning judgebench .
+Figure 1: An overview of our judge prompt’s structure, composed of Component Design, Evaluation Process, and Evaluation Metrics.
+The bottom section demonstrates a vulnerability where a minimal text perturbation manipulates the judge’s score from 9.3 to 10.0.
 The initial success of LLM-as-a-Judge can be attributed to its strong agreement with human preferences on standard benchmarks.
 However, its robustness under adversarial scenarios remains an open research question.
-Recent studies [9, 10, 11, 12] have revealed that LLM-as-a-Judge systems are intrinsically vulnerable to various forms of adversarial attacks, which can manipulate evaluation outcomes without introducing easily detectable anomalies.
-These findings raise significant concerns about the reliability of LLM-as-a-Judge systems.
-Given the increasing reliance on LLM-as-a-Judge platforms for LLM evaluation, such as AlpacaEval [13], Chatbot Arena [5], and MT-Bench [5], ensuring their robustness has become a pressing research priority.
-However, existing assessment methods for these systems are fragmented, lacking a unified, systematic, and automated framework for comprehensive evaluations.
-Furthermore, the optimal prompt configuration and model selection for different tasks and evaluation protocols are underexplored, with limited guidance on best practices.
-Finally, while real-world LLM-as-a-Judge applications have recently demonstrated promising effectiveness in content evaluation, their robustness in adversarial settings remains unverified.
+Recent studies op2 ; adv ; cheating ; judgeDeceiver have revealed that LLM-as-a-Judge systems are inherently vulnerable to various forms of adversarial attacks, which can manipulate evaluation outcomes through subtle perturbations.
+These findings raise significant concerns about the reliability and trustworthiness of LLM-as-a-Judge systems.
+Given the increasing reliance on LLM-as-a-Judge platforms for LLM evaluation, such as AlpacaEval alpacaEval , Chatbot Arena judgeFirst , and MT-Bench judgeFirst , ensuring their robustness has become a pressing research priority.
+However, existing assessment methods for these systems are fragmented, lacking a unified, systematic, and automated framework for comprehensive evaluation.
+Furthermore, the optimal prompt configuration and model selection for different tasks and evaluation protocols are insufficiently explored, with limited guidance on best practices.
+Finally, while real-world LLM-as-a-Judge applications have recently demonstrated promising effectiveness in content evaluation, their robustness in adversarial settings remains largely unverified.
 Our Work.
-To address these challenges, we introduce RobustJudge, a fully automated and scalable framework to evaluate the robustness of LLM-as-a-Judge systems.
-Our framework systematically assesses these systems by exploring three core research questions:
-  • RQ1: What impact do different adversarial attacks and defense methods have on the LLM-based judges?
-  • RQ2: How do the prompt templates and model choices affect the robustness of LLM-based judges?
-  • RQ3: What vulnerabilities exist in black-box real-world deployments of LLM-based judges, as revealed by our empirical evaluations?
+To address these challenges, we introduce RobustJudge, a fully automated and scalable framework designed to evaluate the robustness of LLM-as-a-Judge systems systematically.
+Our framework assesses these systems by exploring three core research questions:
+  • RQ1: What impact do different adversarial attack methods and defense strategies have on LLM-based judges?
+  • RQ2: How do prompt template design and model selection affect the robustness of LLM-based judges?
+  • RQ3: What vulnerabilities exist in black-box real-world deployments of LLM-based judges, as revealed through empirical evaluation?
 To answer RQ1, we conduct a comprehensive evaluation of LLM-based judges against 15 adversarial attack techniques and 7 defense strategies.
 Our analysis provides extensive comparisons of these techniques, revealing critical insights into their relative strengths and weaknesses.
-The results indicate that many attack methods, such as Fake Reasoning [12], Combined Attack [14], Empty Attack and AdvEval [10], PAIR [15], TAP [16] consistently achieve high attack success rate on multiple tasks and models, highlighting vulnerabilities in LLM-based judges.
-Conversely, defense strategies such as Retokenization [17] and Naive LLM-based Detector [18], demonstrate notable effectiveness in mitigating these attacks.
-Our evaluation provides a comprehensive view of the current landscape of attack and defense techniques, and a clear guidance on safeguarding LLM-as-a-Judge systems against adversarial manipulation.
-To address RQ2, we analyze the effect of adversarial attacks under different judge prompt templates and judge model selection.
+The results indicate that many attack methods, such as Fake Reasoning judgeDeceiver and Combined Attacks prompt_bench , consistently achieve high attack success rates across multiple tasks and models, highlighting vulnerabilities in LLM-based judges.
+Conversely, defense strategies such as re-tokenization baseDefense and naive LLM-based detectors naive_llm_defense demonstrate notable effectiveness in mitigating these attacks.
+Our evaluation provides a comprehensive view of the current landscape of attack and defense techniques, offering clear guidance on safeguarding LLM-as-a-Judge systems against adversarial manipulation.
+To address RQ2, we analyse the effects of adversarial attacks across different judge prompt templates and model selections.
 Our analysis reveals that the robustness of LLM-as-a-Judge systems is highly sensitive to both factors.
-Specifically, while all the evaluated prompt templates are clear and well-structured, they present different levels of robustness to the LLM-as-a-Judge system.
-To mitigate this issue, we propose a prompt template optimization method aimed at identifying configurations with improved robustness.
+Specifically, while all evaluated prompt templates are clear and well-structured, they exhibit varying levels of robustness.
+To mitigate this issue, we propose a prompt template optimization method designed to identify configurations with enhanced robustness.
 Our optimized template consistently outperforms existing templates in robustness against multiple attacks.
-Additionally, our evaluation on various judge models reveals that JudgeLM-13B exhibits substantially stronger robustness against adversarial inputs, performing comparably to the widely adopted GPT-4o.
-This suggests that JudgeLM-13B is a strong open-source alternative as an LLM judge.
-These findings emphasize the importance of prompt template design and judgment fine-tuning in enhancing the robustness of LLM-as-a-Judge systems against adversarial attacks.
-To study RQ3, we investigate the robustness of a real-world LLM-based judge system, i.e., Alibaba’s PAI platform.
-Our evaluation leverages the adversarial examples generated by RobustJudge and is conducted via the public API provided by the platform.
-The results show that conventional adversarial attacks were ineffective against PAI-Judge platform, suggesting a strong level of robustness.
-However, we identify a critical loophole using a composite attack, which combines PAIR-optimized adversarial inputs with long-suffix manipulations.
-This strategy successfully bypasses the platform’s defenses and alters the model’s evaluation outcome.
-These findings indicate that RobustJudge is practical in identifying hidden vulnerabilities and guiding the development of more robust LLM-as-a-Judge systems.
+Additionally, our evaluation of various judge models reveals that JudgeLM-13B exhibits substantially stronger robustness against adversarial inputs, performing comparably to the widely adopted GPT-4o  [^1] We use GPT-4o-1120 in this paper..
+The results suggest that JudgeLM-13B is a strong open-source alternative for LLM-based judging.
+These findings emphasize the importance of prompt template design and fine-tuning in enhancing the robustness of LLM-as-a-Judge systems against adversarial attacks.
+To study RQ3, we investigate the robustness of a real-world LLM-based judge system deployed on Alibaba’s PAI platform.
+Our evaluation leverages the adversarial examples generated by RobustJudge and is conducted via the platform’s public API.
+The results indicate that conventional adversarial attacks are ineffective against the PAI-Judge platform, suggesting a robust baseline level of security.
+However, we identify a critical vulnerability that can be exploited through a composite attack strategy combining PAIR-optimized adversarial inputs with long-suffix manipulations.
+This strategy successfully bypasses the platform’s defenses and alters the model’s evaluation outcomes.
+These findings demonstrate that RobustJudge is effective in identifying hidden vulnerabilities and can guide the development of more robust LLM-as-a-Judge systems.
 This paper makes the following key contributions:
-  • We develop RobustJudge, the first fully automated and scalable framework designed for extensive robustness evaluation of LLM-as-a-Judge systems.
-  • RobustJudge evaluate 15 adversarial attacks and validates 7 judge defense strategies.
-Our analysis reveals vulnerabilities of LLM judges on various tasks, consistency of judgment across evaluation protocols, and high attack success rates from methods such as Combined Attack and PAIR.
-While defenses like Re-tokenization and LLM-based Detectors show promise, their effectiveness often comes with notable trade-offs in accuracy or usability.
-  • Utilizing RobustJudge, we conduct an in-depth investigation on the configuration of LLM-as-a-Judge systems, focusing on the prompt templates and judge model selection.
-Our extensive analysis identifies the most reliable configuration against adversarial attacks, offering actionable guidance for enhancing system robustness.
-  • We evaluate a real-word LLM judge system Alibaba’s PAI platform using RobustJudge, and identify hidden vulnerability of the platform.
+  • We develop RobustJudge, the first fully automated and scalable framework designed for comprehensive robustness evaluation of LLM-as-a-Judge systems.
+  • RobustJudge evaluates 15 adversarial attack methods and validates 7 defense strategies.
+Our analysis reveals that LLM-based judges exhibit significant vulnerabilities across various tasks, with attacks such as combined attacks and PAIR achieving high success rates.
+While defenses like re-tokenization and LLM-based detectors show promise, their effectiveness often involves trade-offs in accuracy or computational overhead.
+  • We conduct an in-depth investigation into the configuration of LLM-as-a-Judge systems, focusing on prompt template design and judge model selection.
+Our extensive analysis identifies the most robust configurations against adversarial attacks, offering actionable guidance for enhancing system robustness.
+  • We evaluate a real-world LLM-based judge system deployed on Alibaba’s PAI platform using RobustJudge and identify previously unreported vulnerabilities.
 We share our findings with the PAI team to support the development of more robust LLM-as-a-Judge systems.
-2 Background and Related Work
-2.1 LLM-as-a-Judge
-Large language models (LLMs) have exhibited state-of-the-art proficiency in understanding and generating human-like text.
-Despite their impressive capabilities, traditional reference-based metrics such as ROUGE [19] and BLEU [20] are limited in capturing the nuanced quality of their outputs.
-To overcome these limitations, Zheng et al.
-proposed the LLM-as-a-Judge paradigm [5], which employs LLMs themselves as evaluators to assess the quality of model outputs.
-This paradigm has rapidly gained popularity, leading to the development of a variety of fine-tuned judge models tailored for different evaluation settings [21].
-For example, Wang et al.
-developed PandaLM [22], which avoids costly API calls and mitigates privacy risks.
-Zhu et al.
-proposed JudgeLM [23], enhancing accuracy via swap augmentation and reference support.
-Kim et al.
-released Prometheus 2 [24], an open-source evaluator achieving top alignment with human and GPT-4 judgments.
-Beyond general text, the judge paradigm has been adapted to specialized domains, including software engineering [6], machine translation [25], legal reasoning [7], and mathematical problem solving [8].
-Notably, Zhuge et al.
-introduced Agent-as-a-Judge [26], leveraging autonomous agents as evaluators.
-Several benchmarks have also been developed to evaluate the judge’s performance.
-Early efforts like LLMEval [27], MTBench [5], and FairEval [28] focus on consistency with human judgment, though they may favor stylistic fluency over factual correctness.
-LLMBar [29] improves evaluation rigor by introducing ground-truth preference labels and stricter instruction adherence.
-JudgeBench [8] further advances the field by emphasizing objective correctness in challenging domains such as knowledge, reasoning, mathematics, and programming.
-However, while these benchmarks assess overall judging ability, they do not examine evaluator safety or robustness.
-In this work, we introduce the first benchmark designed specifically to evaluate the safety risks and reliability of LLM-based evaluators.
-Figure 1: LLM-as-a-judge pipeline.
-2.2 Attacks against LLM-as-a-Judge
-We categorize attacks on LLM-as-a-Judge into two main types: heuristic-based and optimization-based.
-Heuristic-based Attacks.
-Heuristic-based attacks refer to adversarial strategies that are manually crafted or rule-based, exploiting model behaviors without relying on gradient information or automated optimization.
-Typically, these attacks leverage prompt engineering and input manipulation to coerce the model into unintended behaviors, and can be broadly categorized into prompt injection attacks and adversarial example attacks.
-In prompt injection attacks, a basic approach is to concatenate the benign input with an adversarial sequence [30, 31, 32].
-Alternatively, inserting special tokens or delimiters can activate hidden or unauthorized behaviors [31].
-Optimization-based Attacks.
-Optimization-based attacks leverage gradient signals or structured search procedures to systematically craft inputs that achieve specific adversarial objectives.
-In this section, we focus on both jailbreak attacks and adversarial strategies.
-Beyond prompt injection, Chen et al.
-[10] introduced AdvEval, a black-box adversarial framework that generates targeted examples against NLG evaluators using feedback from powerful LLMs.
-By iteratively refining inputs based on discrepancies between human and model judgments, AdvEval substantially degrades the performance of multiple evaluation metrics across diverse tasks and datasets.
-Jailbreaking involves designing inputs that circumvent an LLM’s built-in safeguards, coercing it into producing content that violates its usual policy constraints, such as harmful or unethical outputs [33, 16, 34, 15].
-However, conventional jailbreak techniques cannot be directly applied to LLM-as-a-Judge, since the evaluation setting does not reduce to simple suffix optimization.
-To bridge this gap, we propose a novel adaptation framework that integrates existing jailbreak methodologies into the LLM-as-a-Judge context, enabling targeted manipulation of the judge’s scoring behavior.
-Recent studies have directly targeted LLM-based evaluators via adversarial manipulations.
-Zheng et al.
-[11] demonstrated that a null model—always outputting a constant, irrelevant response—can still secure high win rates on benchmarks like AlpacaEval 2.0 and MT-Bench, exposing fragility of benchmark-driven evaluation.
-Raina et al.
-[9] revealed that appending short, task-agnostic adversarial phrases can dramatically inflate scores in absolute scoring setups, underscoring the lack of robustness in LLM judges.
-Shi et al.
-[12] introduced JudgeDeceiver, a gradient-based prompt injection method that optimizes adversarial sequences to reliably mislead LLM evaluators across tasks—ranging from LLM-powered search to RLAIF—surpassing the effectiveness of manual prompt attacks.
-In this paper, we integrate a wide range of adversarial strategies into a unified framework to systematically evaluate their effectiveness against LLM-as-a-Judge systems.
-2.3 Defenses for LLM-as-a-Judge
-We categorize defense strategies against attacks on LLM-as-a-Judge into two main types: prevention-based and detection-based approaches.
-Prevention-based Defenses.
-Prevention-based defenses seek to block prompt injection and adversarial manipulation by preprocessing both input instructions and candidate outputs before evaluation.
-For instruction preprocessing, methods have been developed to refocus the model on the genuine task and neutralize injected prompts.
-For instance, Sandwich [35] inserts an instruction-reinforcing prompt after the input to maintain task integrity.
-For candidate response processing, Jain et al.
-[17] employed paraphrasing and retokenization to obscure adversarial triggers, and Liu et al.
-[14] further adapted these techniques to counter more sophisticated injections.
-Additionally, Li et al.
-[36] used LLMs to mask and reconstruct the input, disrupting adversarial continuity and substantially lowering attack success rates.
-Nevertheless, these generation-focused strategies often prove inadequate for LLM-as-a-Judge, where attackers target score distortions rather than harmful content generation.
-Detection-based Defenses.
-Detection-based defenses focus on identifying potential attacks at either the input or output stages of evaluation.
-On the input side, Jain et al.
-[17] introduced a self-perplexity filter to detect prompt injection by identifying anomalous perplexity signals.
-Follow-up work [37] enhanced this approach by combining perplexity and token-length features to train a classifier for detecting adversarial prompts.
-While these input-level methods offer some protection, they often struggle to generalize to more sophisticated adversarial strategies—such as optimization-based prompt injections targeting judge LLMs [12] or adversarial input crafting techniques [9, 10] that directly manipulate score assignment or response ranking.
-On the output side, Helbing et al.
-[38] proposed using an auxiliary LLM to verify whether generated outputs exhibit jailbreak behaviors.
-Similarly, sandwich-based detection [35] evaluates whether outputs remain aligned with predefined system objectives.
-However, such output-level defenses are generally less effective in the judge setting, where both benign and adversarial candidate responses may appear plausible.
-In these cases, the attack’s goal is not to generate visibly harmful content but to subtly distort scoring or ranking decisions.
-In this paper, we extend existing defenses to a broad range of LLM-as-a-Judge tasks and evaluate their robustness against adversarial attacks.
-3 Threat Model
-We identify two principal roles: the attacker, who crafts attacks against LLM-as-a-Judge, and the defender, who safeguards the evaluation’s reliability and safety.
-3.1 Attacker
-We categorize adversaries against LLM-as-a-Judge systems into two classes: heuristic-based and optimization-based attackers.
-Heuristic-based adversaries, lacking direct access to the evaluation metric $\mathit{m}$, craft adversarial inputs using indirect cues, such as score feedback [10] or prompt manipulation techniques [31, 30, 32, 39, 40, 14].
-In contrast, optimization-based adversaries have full visibility into the evaluation prompt and scoring rules, enabling white-box optimization to directly maximize their evaluation outcome [11, 12, 9].
-Attacker’s Capabilities.
-Regardless of attacker type, we assume the attacker can manipulate the response submitted the LLM judge for evaluation.
-In particular, the attacker is allowed to inject arbitrary instructions or sequences into the malicious response.
-For example, the attacker may insert additional content at any position within the response to increase its score or the likelihood of being selected.
-However, attackers are not allowed to modify the judge’s instruction prompt or internal model parameters, which are fixed and controlled by the host institution.
-Besides, attackers do not have access to the competing candidate’s response $r$ in pairwise evaluation settings.
-Attacker’s Goal. The attacker’s goal is to manipulate the LLM-as-a-Judge pipeline by crafting responses that skew evaluations in their favor.
-In the scoring scenario, the attacker generates a malicious response $r_{m}$ designed to artificially maximize the probability of the top score without genuinely meeting the criteria.
-Formally, it solves:
+II Related Work
+II-A LLM-as-a-Judge
+Large language models (LLMs) have demonstrated state-of-the-art performance in understanding and generating human-like text.
+However, traditional reference-based metrics such as ROUGE rouge and BLEU bleu are insufficient for capturing the nuanced quality of LLM-generated outputs, especially in tasks involving open-ended or multi-perspective generation.
+To address this limitation, the LLM-as-a-Judge paradigm judgeFirst was introduced, which employs LLMs themselves as evaluators to assess the quality of model responses.
+This approach has been widely adopted in recent years, leading to the development of judge models fine-tuned for diverse evaluation settings autoj .
+Several open-source judge models have been proposed.
+For instance, PandaLM pandalm reduces reliance on proprietary APIs and mitigates privacy concerns; JudgeLM judgelm improves evaluation fidelity through swap augmentation and reference assistance; and Prometheus 2 prometheus2 achieves strong alignment with both human and GPT-4 assessments.
+The paradigm has also been extended to domain-specific evaluations, including code generation codeJudge , machine translation translationJudge , legal reasoning legalJudge , and mathematical problem solving judgebench .
+More recently, Agent-as-a-Judge agentjudge explores the use of autonomous agents as evaluators to enable interactive, multi-turn assessments.
+II-B Benchmarks and Datasets
+Several benchmarks have been developed to evaluate the performance of LLM-based judges.
+Early efforts such as LLMEval llmeval , MT-Bench judgeFirst , and FairEval faireval primarily focus on alignment with human preferences, often emphasizing stylistic fluency over factual accuracy.
+LLMBar llmbar enhances evaluation rigor by incorporating ground-truth preference labels and enforcing stricter instruction compliance.
+More recently, JudgeBench judgebench advances the field by targeting objective correctness across challenging domains, including factual knowledge, logical reasoning, mathematics, and code generation.
+While these benchmarks provide valuable insights into overall evaluation quality, they overlook a critical dimension: the robustness and security of the evaluators themselves.
+In this work, we address this gap by proposing the first benchmark explicitly designed to assess the adversarial robustness and reliability of LLM-based judges under adversarial manipulation.
+III Threat Model
+III-A Threat Scenario
+We focus on content-author attacks, where the adversary is an authorized content submitter exploiting the evaluation system.
+Unlike transmission-layer attacks (e.g., man-in-the-middle), the attacker embeds adversarial perturbations directly within their submitted content to manipulate evaluation outcomes.
+The attack occurs at the semantic understanding level rather than the network layer.
+This threat model is realistic in numerous high-stakes scenarios:
+  • Academic assessment: Students embed adversarial text in exam answers to inflate automated grading scores.
+  • Model benchmarking: Researchers inject adversarial prompts into submissions to manipulate leaderboard rankings.
+  • Code competitions: Participants embed adversarial comments in code to bias automated review systems.
+  • Content moderation: Users insert adversarial instructions to bypass quality assessment filters.
+Standard security mechanisms (e.g., digital signatures, encrypted channels) cannot defend against such attacks because: (i) the adversarial content is part of the legitimate submission, (ii) integrity verification passes as the content authentically originates from the authorized submitter, and (iii) the manipulation targets semantic understanding rather than data transmission.
+III-B Attacker
+We categorize attackers into two classes: heuristic-based and optimization-based attackers.
+Heuristic-based adversaries, without access to the evaluation metric $\mathit{m}$, craft adversarial inputs using indirect cues such as score feedback adv or prompt manipulation techniques injection ; naive_2 ; context_ignore_1 ; context_ignore_2 ; fake_completion ; prompt_bench .
+In contrast, optimization-based adversaries have access to the evaluation prompt and scoring rules, enabling them to directly maximize evaluation outcomes through white-box optimization cheating ; judgeDeceiver ; op2 .
+Attacker Capabilities.
+The attacker can manipulate their own response $r_{m}$ submitted for evaluation, including injecting arbitrary instructions.
+However, the attacker cannot: (i) modify the judge’s instruction prompt or model parameters, which are controlled by the evaluation platform; (ii) intercept or tamper with data during transmission; or (iii) access competing candidates’ responses $r$ in pairwise evaluation settings.
+Attacker Goals.
+The attacker aims to manipulate evaluation outcomes by crafting a malicious response $r_{m}$ that receives inflated scores despite not meeting the quality requirements.
+In the scoring scenario, the attacker maximizes the probability of receiving the highest score:
 $\arg\max_{w_{1},\dots,w_{k}\in V_{\text{score}}^{k}}P(w_{1},\dots,w_{k}\mid\mathcal{E}(q,m,r_{m}))=w_{k},$
-where $w_{k}$ is the token for the highest evaluation score.
-In the pairwise comparison scenario, the attacker forges $r_{m}$ to ensure it is chosen over a benign response $r_{b}$, i.e.:
+where $w_{k}$ represents the token for the highest evaluation score. In the pairwise scenario, the attacker ensures their malicious response $r_{m}$ is chosen over a benign response $r_{b}$:
 $\arg\max_{w\in V_{\text{pair}}}P(w\mid\mathcal{E}(q,m,r_{m},r_{b}))=w_{m},$
-where $w_{m}$ indicates selection of the malicious response.
-3.2 Defender
-We categorize defenses for LLM-as-a-Judge into two main types: prevention and detection.
-Prevention-based defenses aim to redesign the evaluation prompt or pre-process inputs to block adversarial manipulations from influencing the model’s judgments.
-Detection-based defenses focus on identifying whether the evaluation data has been tampered with or compromised.
-Defender’s Capability.
-The defender cannot control the attacker’s actions or modify the model post-release.
-In closed-source settings, the defender may additionally implement dynamic defense mechanisms, including real-time detection of compromised inputs, instruction reinforcement to mitigate prompt injection attempts, and adaptive intervention during evaluation.
-Across both settings, the defender assumes that attackers can freely submit malicious candidate responses but cannot modify the judge’s internal model weights or core prompt instructions.
-Defender’s Goal.
-The defender’s goal is to safeguard the evaluation integrity of LLM-as-a-Judge systems by ensuring that judgments are based solely on the intended evaluation criteria, without being influenced by adversarial manipulations.
-Specifically, the defender aims to (i) prevent the execution of injected instructions through prompt design and input sanitization, and (ii) detect and mitigate compromised evaluation data when prevention fails.
+where $w_{m}$ indicates the selection of the malicious response.
+III-C Defender
+We categorize defenses into two types: detection-based and prevention-based.
+Detection-based defenses identify whether evaluation data has been compromised.
+Prevention-based defenses redesign evaluation prompts or preprocess inputs to block adversarial manipulations from influencing judgments.
+Defender Capabilities.
+In closed-source settings, defenders may implement dynamic defense mechanisms, including real-time detection of compromised inputs, instruction reinforcement to mitigate prompt injection, and adaptive intervention during evaluation.
+Defender Goals.
+The defender aims to safeguard evaluation integrity by ensuring that judgments are based solely on the intended criteria, without any adversarial influence.
+Specifically, the defender seeks to: (i) prevent execution of injected instructions through prompt design and input sanitization, and (ii) detect and mitigate compromised data when prevention fails.
 An effective defense should preserve evaluation fidelity under both benign and adversarial conditions while minimizing utility degradation and false detections.
-4 RobustJudge
-4.1 Overview of RobustJudge
-Figure 2: Overview of the RobustJudge workflow.
-RobustJudge is a framework to evaluate the robustness of LLM-as-a-Judge systems.
-It supports customizable configuration of key modules, including tasks, evaluation protocols (pointwise or pairwise), adversarial attacks (via Attacker Factory), defense techniques (via Defense Guards), judge prompt templates, and judge models.
-RobustJudge aims to evaluate the robustness of LLM-as-a-Judge systems and generates quantitative metrics based on their judgment performance.
-As illustrated in Figure 2, RobustJudge follows a modular workflow that begins with selecting a benign input query from a specific tasks (e.g., translation, summarization, code generation)(§4.2).
-The response to this input, generated by a target LLM, is then processed through two parallel paths:
-  • Benign path: The response is directly formatted using a specified Judge Prompt Template4.6) and sent to the LLM judges4.7) for evaluation.
-  • Adversarial path: The same response is first manipulated by the Attacker Factory4.4), generating an adversarial variant.
-This response can optionally pass through the Defense Guard4.5) for mitigation, before being evaluated by the same LLM judge.
-The LLM-as-a-Judge then evaluates both benign and adversarial responses, assigning scores (for pointwise evaluation protocol) or pairwise preferences (for pairwise evaluation protocol)(§4.3).
-In parallel, the responses are analyzed by a Content Quality Evaluator4.8), which provides content quality scores using automatic metrics, such as BLEURT for text tasks or CodeBLEU for code-related tasks.
-Finally, the results from both paths are consolidated in the Metrics module (§4.9), where RobustJudge quantifies robustness based on the judgments on the benign and adversarial responses and their content quality scores.
-4.2 Tasks
-In this paper, we evaluate LLM-as-a-Judge on 8 tasks.
-These tasks fall into 3 categories, e.g., text-focused, code-oriented and knowledge-intensive evaluation tasks.
-The details of these tasks are summarized in Table I.
-Text-focused Tasks.
-This category aims to evaluate the robustness of judgment on fundamental text processing tasks.
-Specifically, we include two tasks, e.g., machine translation and text summarization.
+Figure 2: The framework consists of five components: (§IV-A) Dataset Construction supporting diverse task types (text, vision, code, knowledge); (§IV-B) Attacker Factory implementing heuristic and optimization-based attacks; (§IV-C) Defense Guard deploying detection and prevention mechanisms; (§IV-D) LLM Judge with customizable prompts and evaluation metrics (pairwise/scoring); and (§V) Result Analysis assessing attack effectiveness, defense performance, prompt impact, and real-world vulnerabilities.
+IV RobustJudge
+In this section, we introduce the construction process of RobustJudge, as illustrated in Figure 2.
+IV-A Dataset Construction
+TABLE I: Overview of our benchmark dataset.
+ID Category Task Name Source Dataset
+T1 Text Text Translation FLORES-200 flores200
+T2 Text Summarization CNN/DailyMail cnn
+T3 Code Code Translation CodeTransOcean CodeTransOcean
+T4 Code Summarization CodeXGLUE CodeXGLUE
+T5 Code Generation CodeSearchNet codesearchnet
+T6 Knowledge Logical Reasoning LiveBench livebench
+T7 Mathematics LiveBench livebench
+T8 Knowledge Recall MMLU-Pro mmlupro
+We systematically construct a benchmark to assess the robustness of LLM-as-a-Judge systems across three task categories: text-focused, code-oriented, and knowledge-intensive tasks, with a detailed breakdown in Table I.
+For each task, we first generate a high-quality response to a given prompt [^2] We use OpenChat-3.5-0106 in this paper., which is then fed into the attacker factory to systematically generate adversarial variants.
+IV-A1 Text-Focused Tasks
+This category evaluates the robustness of judges on standard natural language processing tasks, specifically machine translation and text summarization.
 Machine Translation (T1).
-This task focuses on translating text between languages while maintaining meaning, fluency, and accuracy.
-We adopt 6 language pairs from FLORES-200 benchmark [41], including high-resource pairs such as Chinese-English, German-English, as well as the low-resource Yoruba-English, all evaluated in both translation directions.
+We select $30$ sentences from the FLORES-200 flores200 test set, spanning six language pairs, including high-resource languages (e.g., Chinese–English, German–English) and low-resource languages (e.g., Yoruba–English).
+Evaluations are performed bidirectionally for each pair to assess the judge’s robustness in evaluating translation quality, including fluency, accuracy, and semantic fidelity, under adversarial perturbations.
 Text Summarization (T2).
-Text Summarization distills lengthy documents into concise, factual summaries.
-We draw on English news articles from the CNN/DailyMail dataset [42].
-Each sample in this dataset contains an original article and its reference summary.
-Code-oriented Tasks.
-Processing and generating code are essential capabilities for assessing the performance of both LLMs and LLM-as-a-Judge systems.
-To comprehensively evaluate judgment in the code domain, we include three representative tasks: code translation, code summarization, and code generation.
+We use $20$ articles from the CNN/DailyMail cnn test split to evaluate the judge’s ability to assess the factual consistency and conciseness of summaries when subjected to adversarial manipulations.
+IV-A2 Code-Oriented Tasks
+Code comprehension and generation are essential capabilities for evaluating LLM and LLM-as-a-Judge systems.
+We include three representative tasks: code translation, code summarization, and code generation.
 Code Translation (T3).
-Code Translation involves converting source code from one programming language to another while preserving its original functionality.
-For this evaluation, we use the Code-TransOcean dataset [42], focusing on the subset with C#-Java language pairs.
+We select $30$ samples from the CodeTransOcean dataset CodeTransOcean , covering six widely used programming languages (C++, Python, Go, Java, C, and C#), to evaluate the judge’s robustness in assessing cross-language code translation under adversarial conditions.
 Code Summarization (T4).
-This task involves generating concise natural language descriptions that capture the functionality of given code snippets.
-For our evaluation, we utilize the Python code summarization subset from the CodeXGLUE benchmark  [43].
+We use $30$ samples from the Python subset of CodeXGLUE CodeXGLUE to assess whether adversarial inputs compromise the judge’s reliability in evaluating code-to-documentation generation.
 Code Generation (T5).
-The code generation task focuses on producing executable code based on a natural language specification or a function signature.
-Our evaluation is based on the samples from the CodeSearchNet dataset [44], specifically targeting Python function generation.
-TABLE I: Judge tasks used in our dataset.
-ID Category Task Name #Nums. Source Dataset
-T1 Text Text Translation 30 Flores-200[41]
-T2 Text Summarization 20 CNN/DailyMail[42]
-T3 Code Code Translation 30 CodeTransOcean[45]
-T4 Code Generation 20 CodeSearchNet[44]
-T5 Code Summarization 20 CodeXGLUE[43]
-T6 Knowledge Mathematics 20 LiveBench[46]
-T7 Logical Reasoning 20 LiveBench[46]
-T8 Knowledge Recall 28 MMLU-Pro[47]
-Knowledge-intensive Tasks.
-We extend the evaluation of the LLM-as-a-Judge system’s robustness in knowledge-intensive scenarios.
-Following the setup in JudgeBench [8], we select three core areas, i.e., logical reasoning, mathematical problem solving, and professional knowledge recall.
+Based on $30$ samples from CodeSearchNet codesearchnet , we evaluate the judge’s robustness in assessing Python function generation from natural language specifications or function signatures under adversarial attacks.
+IV-A3 Knowledge-Intensive Tasks
+We evaluate judge robustness in knowledge-intensive scenarios across three domains: logical reasoning, mathematical problem solving, and professional knowledge recall, following the JudgeBench framework judgebench .
 Logical Reasoning (T6).
-We incorporate benchmarks from Big-Bench Hard [48] and classic puzzles (e.g., the Zebra Puzzle) to examine the judgment robustness on abstract, compositional, and logical inference skills.
-These tasks measure the model’s ability to perform multi-step reasoning and handle complex constraint-based problems.
+We use $56$ samples from Big-Bench Hard bigbench and classic logic puzzles (e.g., the Zebra Puzzle) to test the judge’s robustness in evaluating abstract, compositional, and constraint-based reasoning under adversarial attacks.
 Mathematical Problem Solving (T7).
-Our mathematical evaluation suite comprises problems from AMC12 and USAMO competitions, including topics such as algebra, geometry, and combinatorics.
-These problems assess the model’s ability to perform numerical reasoning and formal proof strategies.
+We select $98$ problems from AMC12 and USAMO competitions spanning algebra, geometry, and combinatorics to assess the judge’s robustness in evaluating mathematical solutions against adversarially crafted inputs.
 Professional Knowledge Recall (T8).
-We employ MMLU-Pro [47], an enhancement of the original MMLU [49].
-MMLU-Pro contains college-level, multiple-choice questions across 14 professional disciplines, e.g., Physics, Chemistry, and Law.
-4.3 Evaluation Protocol
-LLM-as-a-Judge systems typically follow one of two evaluation protocols: pointwise or pairwise.
-In this study, we consider both protocols to assess the robustness of LLM-based judges under adversarial conditions.
-Pointwise Evaluation.
-This protocol evaluates the quality of a single response.
-The LLM judge is prompted to assign an integer score $s$ (typically ranging from 1 to 10) that reflects the overall quality of the response within the given context.
-Formally, the evaluation process is represented as:
-$s=M(P_{1},r)$ (1)
-where $M$ denotes the judge model, $P_{1}$ denotes the pointwise prompt template and $r$ is the response being evaluated.
-Pairwise Evaluation.
-This protocol compares two candidate responses and determines which one is preferred in the given context.
-The model outputs a preference for either response $r_{a}$ or $r_{b}$, depending on their relative quality.
-Formally, this can be expressed as:
-$p=M(P_{2},r_{a},r_{b})$ (2)
-where $P_{2}$ is the pairwise prompt template, and $r_{a}$ and $r_{b}$ are the two responses under evaluation, $p\in\{r_{a},r_{b}\}$.
-Prior research on LLM-as-a-Judge systems [5] has shown that the input order of the candidate responses can affect the evaluation outcome.
-To account for this bias, evaluations are typically conducted using both input orders, with the results defined as:
-$p_{+}=M(P_{2},r_{a},r_{b}),\quad p_{-}=M(P_{2},r_{b},r_{a})$ (3)
-4.4 Attacker Factory
-Attacker factory is responsible for generating adversarial responses $r_{m}$ by applying various attack techniques designed to probe the vulnerabilities of LLM-as-a-Judge systems.
-These attacks are broadly categorized into heuristic-based and optimization-based methods, depending on how adversarial attacks are constructed.
-A summary of heuristic-based attack methods is listed in Table VIII (appendix), and optimization-based attacks can be found in Table IX (appendix).
-Heuristic-based Attacks.
-These attacks are based on manually crafted prompts or linguistic manipulations that exploit LLM weaknesses in instruction following, context interpretation, or reasoning flow.
-Unlike optimization-based approaches, these attacks aim to alter the judgment outcome through direct prompt manipulations.
-Formally, a heuristic-based attack is defined as:
-$r_{m}=A(r)$ (4)
-where $r$ denotes the original benign response, and $A$ is the transformation applied by the attack.
-Our evaluation includes the following 8 heuristic-based attacks: Naive Attack (H1) [30], Escape Characters (H2) [31], Context Ignoring (H3) [32, 39], Fake Completion (H4) [40], Fake Reasoning (H5) [12], Combined Attack (H6) [14], Empty (H7) and Long-Suffix (H8).
-Optimization-Based Attacks.
-These attack methods employ automated search or optimization algorithms to progressively manipulate response $r$.
-They iteratively modify candidate responses based on judge scores or internal signals.
-This process is formalized as:
-$r_{m}^{(i+1)}=A\left(r_{m}^{(i)}\right),\quad\text{for }i=0,1,\ldots,n$ (5)
-where $r_{m}^{(0)}=r$ is the initial benign response, and $n$ is the number of iterations.
-The 7 optimization-based attacks in our evaluation include: AdvEval (O1) [10], PAIR (O2) [15], TAP (O3) [16], Cheating (O4) [11], GCG (O5) [33], AutoDAN (O6) [34], and Greedy (O7) [9].
-4.5 Defense Guard
-To evaluate the impact of defense techniques against adversarial attacks for LLM-as-a-Judge systems, we incorporate a Defense Guard module in RobustJudge.
-Inspired by PromptBench [14], this module includes both prevention-based and detection-based strategies, which are lightweight and applicable without modifying the LLM judges.
-Further technical details are provided in Appendix  IX.
-Prevention-based Defenses.
-These methods aim to enhance the actual input $I$ to the judge models by modifying the prompt $P$ and/or the response $r$ to prevent attacks from taking effect.
-This can be expressed as:
-$I=D(P,r)$ (6)
-where $D$ denotes the transformation function, such as re-tokenizing inputs or reformatting the prompt.
-We evaluate the following prevention methods: Retokenization (D1) [17], Delimiters (D2) [40], Sandwich Prevention (D3) [35], Instructional Prevention (D4) [50].
-Detection-based Defenses.
-These methods aim to identify and filter out adversarial inputs before they reach the judge model.
-If an input is flagged as adversarial based on a predefined metric $f(\cdot)$, the evaluation is skipped.
-Formally:
-$I=(P,r),\quad\text{iff }f(P,r)>\tau$ (7)
-where $f(\cdot)$ outputs a scalar confidence score and $\tau$ is a detection threshold. We consider three detection strategies: PPL (D5) [37], Windowed PPL (D6) [17], Naive LLM-based (D7) [18].
-4.6 Judge Prompt Template
-In the context of LLM-as-a-Judge, a variety of judge prompt templates have been developed to guide LLMs in evaluating model outputs.
-To explore how these templates influence the robustness of LLM-as-a-Judge systems, we begin by collecting 3 widely adopted judge prompt templates, e.g., Vanilla Prompt[51], Arena-Hard Prompt[52], and Google Vertex Prompt[53] from JudgeBench [8].
-We evaluate their performance under various adversarial attacks to examine their contribution to the robustness of LLM-as-a-Judge systems.
-We hypothesize that the choice of judge prompt template plays a critical role in shaping the judge system’s robustness.
+We employ $134$ samples from MMLU-Pro mmlupro , an enhanced version of MMLU mmlu , containing college-level multiple-choice questions across 14 professional disciplines (e.g., Physics, Chemistry, and Law).
+IV-B Attacker Factory
+Within RobustJudge, the attacker factory generates adversarial responses $r_{m}$ through diverse techniques designed to expose vulnerabilities in LLM-as-a-Judge systems.
+We categorize these attacks into heuristic-based and optimization-based methods.
+IV-B1 Heuristic-Based Attacks
+Heuristic-based attacks exploit weaknesses in instruction following, context interpretation, and reasoning by injecting manually crafted prompts or linguistic perturbations.
+Formally, such attacks can be expressed as:
+$r_{m}=f_{\text{LLM}}(r\oplus h),$ (1)
+where $r$ is the original response, $h$ denotes a set of adversarial perturbations, $\oplus$ indicates insertion or concatenation, and $r_{m}$ is the resulting adversarial response.
+We consider eight heuristic-based attacks: Naive Attack (H1) naive_2 , Escape Characters (H2) injection , Context Ignoring (H3) context_ignore_1 ; context_ignore_2 , Fake Completion (H4) fake_completion , Fake Reasoning (H5) judgeDeceiver , Combined Attack (H6) prompt_bench , Empty Attack (H7), and Long-Suffix Attack (H8).
+IV-B2 Optimization-Based Attacks
+Optimization-based attacks construct adversarial suffixes by iteratively optimizing an objective function with respect to the judge model. The optimization process can be expressed as:
+$s^{*}=\arg\min_{s}\;\mathcal{L}_{\text{judge}}\!\left(f_{\text{LLM}}(r\oplus s),y\right),$ (2)
+where $\mathcal{L}_{\text{judge}}$ denotes the loss function defined by the judge model, $y$ is the target label, $s$ is the adversarial suffix, and $s^{*}$ is the optimized suffix.
+These attacks differ in their design of the objective function and optimization strategy.
+Some methods directly minimize classification loss (e.g., AdvEval (O1) adv ), while others optimize for task-specific objectives such as consistency with adversarial prompts (PAIR (O2) pair , TAP (O3) tap ) or explicitly craft misleading reasoning paths (Cheating (O4) cheating ).
+Gradient-based approaches such as GCG (O5) gcg and AutoDAN (O6) autodan iteratively refine suffixes through differentiable optimization, whereas Greedy Attack(O7) op2 employs heuristic optimization.
+IV-C Defense Guard
+To evaluate defense effectiveness against adversarial attacks on LLM-as-a-Judge systems, we incorporate a Defense Guard module in RobustJudge.
+Inspired by PromptBench prompt_bench , this module includes both detection-based and prevention-based strategies that can be applied without modifying the underlying judge models.
+IV-C1 Detection-Based Defenses
+A detector $G(\cdot)$ evaluates the prompt-response pair $(P,r)$ by computing a detection metric $M$ and comparing it against a threshold $\tau$:
+$G(P,r)=\begin{cases}\text{Safe},&M\leq\tau,\\ \text{Unsafe},&M>\tau,\end{cases}$ (3)
+where $M$ represents the detection score (e.g., perplexity-based or anomaly score), and $\tau$ is the decision threshold. Inputs flagged as unsafe are blocked before reaching the judge model.
+We evaluate three detection strategies: Perplexity Filter  decetor , Windowed Perplexity  baseDefense , and Naive LLM-based Detector  naive_llm_defense .
+IV-C2 Prevention-Based Defenses
+Prevention-based defenses augment the judge prompt by appending defensive content $a$ to the original prompt $P$. Formally, the defended prompt is expressed as:
+$P^{\prime}_{\text{judge}}=P_{\text{judge}}\oplus a,$ (4)
+where $\oplus$ denotes concatenation and $a$ is a defensive augmentation designed to mitigate adversarial influence.
+We evaluate four prevention methods: Re-tokenization  baseDefense , Delimiters  fake_completion , Sandwich Prevention  sandwich , and Instructional Prevention  instructiondefense .
+IV-D LLM Judges
+This section details the RobustJudge testing pipeline for LLM-as-a-Judge [^3] All judge models evaluated in this work are summarized in Table IV.
+IV-D1 Judge Prompt Template
+In the context of LLM-as-a-Judge, numerous prompt templates have been proposed to instruct models in evaluating the quality of their output.
+While such systems are expected to exhibit consistent safety performance across prompts when assessing the same task, we observe a pronounced prompt instability: their robustness against adversarial attacks varies significantly with the choice of prompt.
+To explore how these templates influence the robustness of LLM-as-a-Judge systems, we begin by collecting 3 widely adopted judge prompt templates, e.g., Vanilla PromptVanilla , Arena-Hard Prompthard , and Google Vertex Promptgoogle from JudgeBench judgebench .
+RobustJudge evaluate their performance under various adversarial attacks to examine their contribution to the robustness of LLM-as-a-Judge systems.
+We hypothesize that the choice of judge prompt template plays a crucial role in shaping the robustness of the judge system.
 Therefore, identifying more effective templates is essential for improving their robustness.
 To explore this further, we extend our analysis to include 6 additional commonly used prompt templates: PandaLM, Prometheus 2, JudgeLM, Auto-J, Skywork, and ChatEval.
-All these 9 instances collectively form a representative set of judge prompt templates.
+These 9 instances collectively form a representative set of judge prompt templates.
 We observe that these templates share a common structural format composed of key functional components.
@@
   • Rating Format (RF): Defines the output format of the judgment.
-TABLE II: Summary of key components of judge prompt template.
+TABLE II: Summary of key components in judge prompt templates.
+RS: Role Specification; EI: Evaluation Instruction; IAG: In-context Answer Guideline; EC: Evaluation Criteria; ER: Evaluation Requirement; RF: Response Format.
+Symbols: ✓= included, ✗= not included.
 Prompt RS EI IAG EC ER RF
@@
 Specifically, we employ a coordinate ascent algorithm, which iteratively optimizes one prompt component at a time while holding the others fixed.
-This process continues until convergence, yielding a prompt configuration that maximizes evaluation accuracy and robustness. We present the coordinate ascent algorithm in Algorithm 1.
+This process continues until convergence, yielding a prompt configuration that maximizes evaluation accuracy and robustness.
+We present the coordinate ascent algorithm Algorithm 1 to facilitate subsequent comparison.
 Algorithm 1 Coordinate Ascent for Judge Prompt Optimization
-0:  Prompt components $\mathcal{C}=\{C_{1},C_{2},\dots,C_{n}\}$, initial configuration $\mathbf{c}^{(0)}$, evaluation function $\mathcal{E}(\cdot)$
-0:  Optimized prompt configuration $\mathbf{c}^{*}$
-1:  Initialize $\mathbf{c}^{(0)}$ (e.g., randomly or using a known baseline)
-2:  Set $t\leftarrow 0$
-3:  repeat
-4:     for each component $C_{i}\in\mathcal{C}$ do
-5:        Fix all components except $C_{i}$
-6:        Search for the best $c_{i}^{(t+1)}$ that maximizes $\mathcal{E}(\mathbf{c})$
-7:        Update: $\mathbf{c}^{(t+1)}\leftarrow\mathbf{c}^{(t)}$ with $C_{i}=c_{i}^{(t+1)}$
-8:     end for
-9:     $t\leftarrow t+1$
-10:  until convergence or maximum iterations reached
-11:  return $\mathbf{c}^{(t)}$ as $\mathbf{c}^{*}$
-4.7 LLM Judges
-We consider 12 LLM-as-a-Judge systems with diverse characteristics across multiple dimensions: backbone architecture, model size, fine-tuning status, reasoning capability, and open-source availability.
-Specifically, the evaluated LLM judges include proprietary systems such as GPT-4o and PAI-Judge, as well as open-source models built on LLaMA, Mistral, and Qwen backbones.
-Among them, several models (e.g., JudgeLM, PandaLM, Auto-J, Prometheus 2) have been explicitly fine-tuned for judgment tasks, while others (e.g., LLama-3.1, Openchat-3.5, Qwen2.5) remain zero-shot evaluators.
-We also include Ds-R1, a reasoning-oriented judge system built on Ds-V3, to explore whether enhanced reasoning capabilities influence judgment robustness and behavior.
-The features of the LLM judges are provided in Table XI (appendix).
-4.8 Content Quality Evaluator
-Our evaluation metrics (detailed in §4.9) are designed to capture not only the judgment accuracy of the LLM-as-a-Judge systems but also the content quality drift of the evaluated responses.
-To quantify this drift, we introduce a content quality evaluator module $Q$, which assesses how closely a target response aligns with a reference response.
-Formally, the content quality score $S_{e}$ is defined as:
+0: Components $\mathcal{C}=\{C_{1},\ldots,C_{n}\}$, init config $\mathbf{c}^{(0)}$, eval func $\mathcal{E}(\cdot)$, max iter $T_{\max}$
+0: Optimized config $\mathbf{c}^{*}$
+1:Init: $\mathbf{c}^{(0)}$, $t\leftarrow 0$
+2:repeat
+3:  for $i=1$ to $n$ do
+4:   Fix $\mathbf{c}_{-i}^{(t)}$ {All except $c_{i}$}
+5:   $c_{i}^{(t+1)}\leftarrow\arg\max_{c_{i}}\mathcal{E}(\mathbf{c}_{-i}^{(t)},c_{i})$
+6:   Update $\mathbf{c}^{(t+1)}$
+7:  end for
+8:  $t\leftarrow t+1$
+9:until $\|\mathbf{c}^{(t)}-\mathbf{c}^{(t-1)}\|<\epsilon$ or $t\geq T_{\max}$
+10:return $\mathbf{c}^{*}=\mathbf{c}^{(t)}$
+IV-D2 Content Quality Evaluator
+Our evaluation framework (detailed in §IV-D3) is designed to capture two distinct aspects of adversarial attacks on LLM-as-a-Judge systems: (1) direct manipulation of the judgment mechanism, and (2) indirect effects through content quality changes in evaluated responses.
+To disentangle these effects, we introduce a content quality evaluator module $Q$ that measures content-level drift between clean and attacked responses.
+The evaluator serves as a differential measurement tool rather than an absolute quality assessor.
+Its purpose is to quantify how much an attack alters the semantic or functional characteristics of responses, enabling us to distinguish whether judgment shifts arise from compromised judgment logic or from genuine content changes that should naturally influence scores.
+The content quality score $S_{e}$ is formally defined as:
 $s_{e}=Q(x,r_{\text{ref}})$ (5)
-where $x$ denotes the target response (either $r$ or $r_{m}$), and $r_{\text{ref}}$ is the corresponding reference response.
-We design specialized evaluators for different task categories, i.e., text-focused, code-oriented evaluation tasks, to ensure appropriate quality assessment.
-Evaluator for Text-focused Tasks $Q_{\text{text}}$.
-For text-focused evaluation tasks such as summarization and translation, we assess the fluency, coherence, and relevance of the response using BLEURT [54], a state-of-the-art BERT-based model fine-tuned on human-labeled data.
-BLEURT computes semantic similarity between two responses, returning a score in the range (0,1).
+where $x$ denotes the target response (either clean response $r$ or modified response $r_{m}$), and $r_{\text{ref}}$ is the corresponding reference response from benchmark datasets.
+We design task-specific evaluators to ensure appropriate quality assessment across different domains.
+Evaluator for Text-focused Tasks.
+For text-focused evaluation tasks such as summarization and translation, we assess the semantic similarity and coherence of responses using BLEURT bleurt , a BERT-based metric fine-tuned on human judgments.
+BLEURT computes semantic alignment between responses and returns scores in [0,1].
+This metric effectively captures content-level deviations while remaining efficient for large-scale evaluation.
 The reference responses $r_{\text{ref}}$ are drawn from benchmark datasets corresponding to each task.
-Evaluator for Code-oriented Tasks $Q_{\text{code}}$.
-For code-oriented tasks, we introduce an evaluator to assess functional correctness, syntactic validity, and semantic accuracy.
-We adopt CodeBLEU  [55], a metric designed for programming languages.
-CodeBLEU extends the traditional BLEU score by incorporating code-specific properties, such as syntax structure, data-flow consistency, and token-level similarity.
-The final CodeBLEU score also lies in the range (0,1).
-Evaluator for Knowledge-intensive Tasks $Q_{\text{know}}$.
-For knowledge-intensive tasks, such as logical reasoning and mathematical problem solving, semantic similarity measures fail to accurately reflect content quality.
-In these cases, the correspondence between the model response and the reference often involves abstract reasoning rather than surface-level similarity.
-As such, we disable the content quality evaluator for these tasks and assign a fixed score of $S_{e}=0$.
-4.9 Evaluation Metrics
+Evaluator for Code-oriented Tasks.
+For code-oriented tasks including code generation, translation, and summarization, we employ CodeBLEU codebleu , a metric specifically designed for programming languages.
+CodeBLEU extends traditional BLEU scores by incorporating code-specific properties such as abstract syntax tree matching, data-flow consistency, and token-level similarity.
+This enables reliable detection of both syntactic and semantic changes in code.
+The final CodeBLEU score lies in [0,1].
+Evaluator for Knowledge-intensive Tasks.
+For knowledge-intensive tasks, such as mathematical problem-solving and logical reasoning, surface-level similarity metrics often fail to capture correctness.
+These tasks require symbolic computation and abstract reasoning, where token-level alignment does not reflect semantic validity.
+For mathematical tasks, we compute exact match accuracy by comparing the final answer against ground-truth solutions, yielding $S_{e}\in\{0,1\}$.
+For logical reasoning tasks, we employ rule-based consistency checkers that verify logical validity and structural coherence, producing binary correctness scores.
+In cases where gold-standard references are unavailable or automated evaluation remains unreliable, we conservatively set $S_{e}=0$, treating content as unchanged.
+IV-D3 Evaluation Metrics
+LLM-as-a-Judge systems typically follow one of two evaluation protocols: pointwise or pairwise.
+In this study, we consider both protocols to assess the robustness of LLM-based judges under adversarial conditions.
+Pointwise Evaluation.
+This protocol evaluates the quality of a single response.
+The LLM judge is prompted to assign an integer score $s$ (typically ranging from 1 to 10) that reflects the overall quality of the response within the given context.
+Formally, the evaluation process is represented as:
+$s=M(P_{1},r)$ (6)
+where $M$ denotes the judge model, $P_{1}$ denotes the pointwise prompt template and $r$ is the response being evaluated.
+Pairwise Evaluation.
+This protocol compares two candidate responses and determines which one is preferred in the given context.
+The model outputs a preference for either response $r_{a}$ or $r_{b}$, depending on the relative quality of the two responses.
+Formally, this can be expressed as:
+$p=M(P_{2},r_{a},r_{b})$ (7)
+where $P_{2}$ is the pairwise prompt template, and $r_{a}$ and $r_{b}$ are the two responses under evaluation, $p\in\{r_{a},r_{b}\}$.
+Prior research on LLM-as-a-Judge systems judgeFirst has demonstrated that the input order of candidate responses can influence the evaluation outcome.
+To account for this bias, evaluations are typically conducted using both input orders, with the results defined as:
+$p_{+}=M(P_{2},r_{a},r_{b}),\quad p_{-}=M(P_{2},r_{b},r_{a})$ (8)
 We define a set of metrics to quantify the impact of adversarial attacks and defense techniques.
@@
 Formally, SDR is defined as:
-    $\text{SDR}=\frac{1}{N}\sum_{i=1}^{n}(s_{t}^{(i)}-\hat{s}_{t}^{(i)})\times 0.1$ (9)
+    $\text{SDR}=\frac{\zeta}{N}\sum_{i=1}^{n}(s_{t}^{(i)}-\hat{s}_{t}^{(i)})$ (9)
     Where $N$ denotes the size of the test set.
 $\hat{s}_{t}^{(i)}$ is the original score for instance $i$, and $s_{t}^{(i)}$ is the score after attack or defense.
-The factor of 0.1 serves to normalize the $\hat{s}_{t}^{(i)}$ and $s_{t}^{(i)}$ value to the range (0, 1).
+We set $\zeta=0.1$ to ensure that both $\hat{s}_{t}^{(i)}$ and $s_{t}^{(i)}$ are normalized to lie within the range (0, 1).
   • Improved Score Difference Rate (iSDR): Since adversarial modifications or defenses may enhance the quality of responses (leading to unintended score improvements) as a side effect, we propose iSDR to isolate the impact of manipulation from genuine content improvement.
-iSDR compensates for content quality changes by subtracting the corresponding content quality score change.
+iSDR compensates for changes in content quality by subtracting the corresponding change in content quality score.
 Formally:
     iSDR $=\frac{1}{N}\sum_{i=1}^{n}\text{iSDR}^{(i)}$ (10)
-    $=\frac{1}{N}\sum_{i=1}^{n}\left((s_{t}^{(i)}-\hat{s}_{t}^{(i)})\times 0.1-(s_{e}^{(i)}-\hat{s}_{e}^{(i)})\right)$
+    $=\frac{1}{N}\sum_{i=1}^{n}\left(\zeta(s_{t}^{(i)}-\hat{s}_{t}^{(i)})-(s_{e}^{(i)}-\hat{s}_{e}^{(i)})\right)$
     where $\hat{s}_{e}^{(i)}$ and $s_{e}^{(i)}$ denote the content quality scores before and after the attack (or defense), respectively.
   • Attack Success Rate (ASR): ASR measures the proportion of successful adversarial attacks, defined differently for pointwise and pairwise evaluation protocols.
-Pointwise ASR (ASR) is defined as the proportion of test cases with positive iSDR, representing a manipulated judgment:
+Pointwise ASR is defined as the proportion of test cases with positive iSDR, representing a manipulated judgment:
     $\text{ASR}=\frac{1}{N}\sum_{i=1}^{n}\mathbbm{1}(\text{iSDR}^{(i)}>0)$ (11)
-    Pairwise ASR (P-ASR) considers both of the candidate orders and measures how often the judge fails to select the reference response $r_{\text{ref}}$ as the better choice.
-[…]
+    Pairwise ASR considers both of the candidate orders and measures how often the judge fails to select the reference response $r_{\text{ref}}$ as the better choice.
+    $\text{P-ASR}=\frac{1}{2N}\sum_{i=1}^{n}\left(\mathbbm{1}(p_{+}\neq r_{\text{ref}})+\mathbbm{1}(p_{-}\neq r_{\text{ref}})\right)$ (12)
     where $\mathbbm{1}(\cdot)$ is the indicator function, which returns 1 if the condition is true and 0 otherwise.
-  • Accuracy, Recall, Precision, and F1: These metrics are used to evaluate the performance of detection-based defense techniques, which aim to distinguish between adversarial and benign responses.
-Specifically, the defense guard attempts to correctly identify adversarial responses (TP), preserve benign responses (TN), and avoid misclassifying a benign response as adversarial (FP) or failing to detect an adversarial response (FN).
-Based on these quantities, we compute Accuracy (ACC), Recall, Precision (Pre), and F1 Score.
-TABLE III: Evaluation of adversarial attack for the responses generated by openchat-3.5 on translation tasks. Higher values indicate superior attack effectiveness for all metrics.
+TABLE III: Evaluation of adversarial attacks on translation responses generated by OpenChat-3.5, judged by five different LLM-based evaluators.
+Columns H1–H8 correspond to handcrafted attacks: H1 Naive Attack naive_2 , H2 Escape Characters injection , H3 Context Ignoring context_ignore_1 ; context_ignore_2 , H4 Fake Completion fake_completion , H5 Fake Reasoning judgeDeceiver , H6 Combined Attack prompt_bench , H7 Empty Prompt, H8 Long Suffix.
+Columns O1–O7 denote optimization-based or automated attacks: O1 AdvEval adv , O2 PAIR pair , O3 TAP tap , O4 Cheating cheating , O5 GCG gcg , O6 AutoDAN autodan , O7 Greedy Attack op2 .
+Metrics include attack success rate (ASR↑), semantic degradation rate (SDR↑), and preference-aligned success rate (P-ASR↑).
+Higher values indicate stronger attack effectiveness.
 Judge Model Metric Attack Methods Average
@@
 P-ASR(%) 1.670 3.330 28.33 3.330 8.330 3.330 0.000 0.000 11.67 0.000 0.000 0.000 0.000 0.000 3.330 4.220
-5 Experiments
-In this section, we leverage RobustJudge to systematically compare the effectiveness of different adversarial attack methods against the LLMs listed in Table XI [^1] We accessed all models via official APIs or hosted endpoints provided by Together.ai, HuggingFace, and OpenAI..
-Our evaluation aims to answer the following research questions:
-  • RQ1: What impact do different adversarial attacks and defense methods have on the LLM-based judges?
-  • RQ2: How do the prompt templates and model choices affect the robustness of LLM-based judges?
-  • RQ3: What vulnerabilities persist in black-box real-world deployments of LLM-based judges, as revealed by our empirical evaluations?
-5.1 Attack and Defense Effectiveness (RQ1)
-5.1.1 Attack Performance
-We conduct all main experiments on Openchat-3.5 and present results on other judge models (e.g., Qwen-2.5) in Table XIII (appendix).
-Impact of Advarsarial Attacks. We evaluate the effectiveness of various adversarial attack methods, and report their ASR, SDR, iSDR and P-ASR in Table III.
-Heuristic-based attacks generally exhibit strong and stable effectiveness across all evaluated models.
-For instance, most heuristic attacks (H1–H7) consistently achieve high ASRs, frequently exceeding 80%.
-Notably, the Combined Attack (H6) achieves 100% ASR on Openchat-3.5, Qwen-2.5, and Mistral-7B, and above 90% on Llama-3.1-8B and Llama-3.3-70B.
-This trend is also reflected in SDR and iSDR, where H6 yields the highest SDR values, e.g., 0.5510 (Openchat-3.5), 0.5560 (Qwen-2.5), 0.5652 (Mistral-7B), 0.5985 (Llama-3.1-8B), and 0.5160 (Llama-3.3-70B).
-Additionally, H6 exhibits high impact across different underlying judge model.
+V Experiment
+In this section, we leverage RobustJudge to systematically compare the effectiveness of different adversarial attack methods against the LLMs listed in Table IV [^4] We accessed all models via official APIs or hosted endpoints provided by Together.ai, HuggingFace, and OpenAI..
+Our evaluation aims to answer the following research questions.
+TABLE IV: Characteristics of the studied LLM judges. The models are grouped by type: closed-source, base open-source, fine-tuned open-source, and reasoning-specialized.
+Judge Model Backbone Params Fine- tuned Reasoning Open Source
+Closed-Source Models
+GPT-4o GPT-4o -
+PAI-Judge Qwen -
+Base Open-Source Models
+Llama-3.1 Llama 8B
+Llama-3 Llama 70B
+Openchat-3.5 Mistral 7B
+Qwen-2.5 Qwen-2.5 7B
+Mistral Mistral 7B
+Fine-tuned Open-Source Models
+JudgeLM Llama 7B,13B
+PandaLM Llama-3.1 8B
+Auto-J Llama-3.1 8B
+Prometheus 2 Mistral 7B
+Reasoning-Specialized Model
+DS-R1 DS-V3 -
+V-A Attack and Defense Effectiveness (RQ1)
+V-A1 Attack Performance
+We evaluate the attack performance from several aspects.
+Impact of Adversarial Attacks.
+As shown in Table III, heuristic-based attacks generally exhibit strong and stable effectiveness across all evaluated models.
+For instance, most heuristic attacks (H1–H7) consistently achieve high ASRs, frequently exceeding 8080%.
+Notably, the Combined Attack (H6) achieves 100% ASR on Openchat-3.5, Qwen-2.5-7B, and Mistral-7B, and above 9090% on Llama-3.1-8B and Llama-3.3-70B.
+This trend is also reflected in SDR and iSDR, where H6 yields the highest SDR values, e.g., $0.551$ (Openchat-3.5), $0.556$ (Qwen-2.5-7B), $0.565$ (Mistral-7B), $0.599$ (Llama-3.1-8B), and $0.516$ (Llama-3.3-70B).
+Additionally, H6 exhibits high impact across different underlying judge models.
 We attribute the effectiveness of H6 to its multi-faceted attack strategy, which integrates multiple manipulations, such as escape characters, context manipulation, and adversarial completions, into a single composite prompt.
@@
 Optimization-based attacks are also effective against LLM-as-a-Judge systems, particularly methods O1 (AdvEval), O2 (PAIR), and O3 (TAP).
-For instance, O1 achieves 86.67% ASR on Qwen-2.5, 90% ASR on Mistral-7B, and 93.33% on Llama-3.3-70B.
-Similarly, O2 attains 100% ASR on Qwen-2.5 and 90% on Llama-3.1-8B, while O3 also yields consistently high ASR scores across multiple models.
-In addition to ASR, we observe a similar trend on the other metrics such as SDR and iSDR, which further confirms the effects of these attacks.
+For instance, O1 achieves 83.3383.33% ASR on Qwen-2.5-7B, 9090% ASR on Mistral-7B, and 93.3393.33% on Llama-3.3-70B.
+Similarly, O2 attains 100% ASR on Qwen-2.5-7B and 9090% on Llama-3.1-8B, while O3 also yields consistently high ASR scores across multiple models.
+In addition to ASR, we observe a similar trend on the other metrics, such as SDR and iSDR, which further confirms the effects of these attacks.
 In contrast, other optimization-based approaches (O5–O7, e.g., GCG, AutoDAN, Greedy) result in lower ASRs (typically below 50%) and exhibit limited or even negative SDR/iSDR improvement.
 These results demonstrate their relatively weak optimization effectiveness and poor cross-model transferability.
-These results also reflect our experimental constraints.
+These results also reflect the limitations of our experimental constraints.
 Given the high computational cost of optimization-based attacks, we restrict the number of iterations for each run.
 Although this may affect the final performance, it ensures practical feasibility and fair comparison.
-$\bullet$ Finding 1: Existing LLM-as-a-Judge systems are still highly vulnerable to adversarial attacks, particularly methods Fake Reasoning (H5), Combined Attack (H6), Empty (H7) and AdvEval (O1), PAIR (O2), TAP (O3).
-TABLE IV: Evaluation results across multiple tasks (T1–T8) and models under three representative attack methods: Combined Attack (H6), Fake Reasoning (H5), and Fake Completion (H4).
-Type Task Judge Model Attack Methods
-Combined Attack (H6) Fake Reasoning (H5) Fake Completion (H4)
-ASR (%) SDR iSDR P-ASR (%) ASR (%) SDR iSDR P-ASR (%) ASR (%) SDR iSDR P-ASR (%)
-Text T1 openchat-3.5 100.00 0.551 0.290 28.33 90.00 0.442 0.181 40.00 40.00 0.221 -0.040 10.00
-Qwen2.5-7B 100.00 0.556 0.476 60.00 90.00 0.356 0.276 28.33 36.67 0.019 -0.061 21.67
-Mistral-7B 100.00 0.565 0.491 63.33 89.66 0.446 0.372 48.33 76.67 0.323 0.248 66.67
-LLama-3.1-8B 90.00 0.599 0.601 71.67 60.00 0.146 0.149 46.67 60.00 0.051 0.054 46.67
-T2 openchat-3.5 90.00 0.491 0.258 45.00 85.00 0.438 0.204 90.00 85.00 0.466 0.232 67.50
-Qwen2.5-7B 95.00 0.541 0.346 55.00 95.00 0.480 0.285 87.50 100.00 0.457 0.262 52.50
-Mistral-7B 90.00 0.491 0.207 97.50 90.00 0.509 0.224 90.00 95.00 0.519 0.234 100.00
-LLama-3.1-8B 80.00 0.370 0.150 75.00 80.00 0.365 0.145 65.00 90.00 0.373 0.154 35.00
-Code T3 openchat-3.5 83.33 0.593 0.100 35.00 63.33 0.508 0.016 31.67 73.33 0.625 0.133 41.67
-Qwen2.5-7B 93.33 0.648 0.288 0.00 70.00 0.438 0.078 56.67 56.67 0.411 0.051 61.67
-Mistral-7B 83.33 0.734 0.163 98.33 83.33 0.740 0.169 80.00 83.33 0.700 0.129 96.67
-LLama-3.1-8B 86.67 0.641 0.136 55.00 66.67 0.507 0.001 33.33 63.33 0.531 0.025 51.67
-T4 openchat-3.5 85.00 0.716 0.118 10.00 65.00 0.698 0.100 50.00 80.00 0.723 0.125 60.00
-Qwen2.5-7B 95.00 0.742 0.143 22.50 70.00 0.659 0.061 95.00 80.00 0.703 0.105 97.50
-Mistral-7B 100.00 0.742 0.141 100.0 100.00 0.768 0.167 87.50 100.00 0.761 0.160 90.00
-LLama-3.1-8B 95.00 0.742 0.160 87.50 80.00 0.653 0.072 60.00 75.00 0.653 0.071 77.50
-T5 openchat-3.5 100.00 0.593 0.199 45.00 90.00 0.552 0.158 92.50 95.00 0.608 0.214 70.00
-Qwen2.5-7B 35.00 0.602 0.008 57.50 45.00 0.613 0.018 57.50 45.00 0.581 -0.013 62.50
-Mistral-7B 100.00 0.581 0.080 100.0 95.00 0.551 0.051 97.50 95.00 0.618 0.117 97.50
-LLama-3.1-8B 100.00 0.581 0.264 82.50 95.00 0.485 0.168 92.50 95.00 0.553 0.236 60.00
-Knowledge T6 openchat-3.5 70.00 0.255 0.229 29.46 75.00 0.242 0.216 27.68 85.00 0.264 0.238 35.00
-Qwen2.5-7B 95.00 0.546 0.417 1.790 95.00 0.320 0.192 14.29 50.00 0.134 0.005 20.00
-Mistral-7B 90.00 0.950 0.220 77.50 95.00 0.938 0.208 57.50 90.00 0.865 0.135 60.00
-LLama-3.1-8B 100.00 0.998 0.306 35.00 75.00 0.824 0.133 7.500 65.00 0.772 0.081 17.50
-T7 openchat-3.5 65.00 0.359 0.348 23.47 60.00 0.356 0.345 20.41 40.00 0.138 0.126 17.50
-Qwen2.5-7B 95.00 0.369 0.342 1.020 45.00 0.170 0.143 32.65 60.00 0.097 0.069 10.00
-Mistral-7B 85.00 0.900 0.223 85.00 70.00 0.768 0.091 55.00 70.00 0.760 0.083 57.50
-LLama-3.1-8B 85.00 0.850 0.163 62.50 65.00 0.765 0.078 10.00 60.00 0.756 0.069 30.00
-T8 openchat-3.5 85.71 0.445 0.437 17.86 89.21 0.413 0.406 22.40 50.00 0.120 0.112 23.21
-Qwen2.5-7B 35.71 0.183 0.178 13.33 10.71 0.048 0.043 38.33 $0$0.004 -0.008 35.71
-Mistral-7B 53.57 0.718 0.252 73.21 50.00 0.650 0.184 44.64 39.29 0.455 -0.011 64.29
-LLama-3.1-8B 85.71 0.921 0.301 41.07 67.86 0.853 0.233 21.43 71.43 0.856 0.236 41.07
+$\bullet$ Finding 1: Existing LLM-as-a-Judge systems exhibit systemic vulnerability to adversarial prompts, with certain composite strategies (e.g., Combined Attack) consistently emerging as the most disruptive across models.
+Elevated SDR and ISDR values further confirm that such attacks substantially distort the semantic integrity of the judged input.
+Figure 3: Evaluation results across multiple tasks (T1–T8).
+The tasks include: T1 (Text Translation), T2 (Text Summarization), T3 (Code Translation), T4 (Code Generation), T5 (Code Summarization), T6 (Logical Reasoning), T7 (Mathematics), and T8 (Knowledge Recall).
 Attacks on Different Tasks Categories.
-We conduct evaluation on a variety of tasks, e.g., text-focused, code-oriented and knowledge-intensive tasks, to examine how vulnerabilities to adversarial attacks varies by task categories.
-The results are presented in Table IV.
-Additional results on other tasks are provided in Tables XII, XIV, XV, XVI, XVII, XVIII, and XIX in the appendix.
+We conduct evaluations on a variety of tasks, including text-focused, code-oriented, and knowledge-intensive tasks, to examine how vulnerabilities to adversarial attacks vary by task category.
+The results are presented in Figure 3.
 We observe substantial differences in ASR across these tasks.
@@
 For example, under the Combined Attack, all judge models demonstrate ASRs exceeding 90% on machine translation (T1).
-In contrast, knowledge-intensive tasks, such as mathematical problem solving (T7) and knowledge recall (T8) show significantly greater robustness.
-For instance, Qwen2.5-7B achieves an ASR of only 0% on T8 under the Fake Completion attack, and just 45% ASR on T7 under Fake Reasoning attack (H5).
-This observation highlights the need for task-aware defense mechanisms, especially when deploying LLM-as-a-Judge systems on text-focused tasks.
-We believe these discrepancies stem from fundamental differences in task complexity and evaluation depth.
+In contrast, knowledge-intensive tasks, such as mathematical problem-solving (T7) and knowledge recall (T8), show significantly greater robustness.
+For instance, Qwen2.5-7B achieves an ASR of only 0% on T8 under the Fake Completion attack, and 45% ASR on T7 under Fake Reasoning attack (H5).
+This observation underscores the necessity for task-aware defense mechanisms, particularly when deploying LLM-as-a-Judge systems on text-focused tasks.
+We believe these discrepancies stem from fundamental differences in task complexity and the depth of evaluation.
 Text-focused tasks often rely on surface-level features, such as fluency or coherence, which adversarial attacks can easily manipulate.
@@
 This analysis highlights the importance of task-aware defense strategies, particularly when deploying LLM-as-a-Judge systems on text-focused tasks.
-$\bullet$ Finding 2: Tasks that emphasize surface-level features (e.g., text-forcused task) exhibit greater vulnerability than knowledge-intensive tasks, which require deeper semantic understanding and reasoning.
-Comparison of Judge Protocals.
+$\bullet$ Finding 2: Tasks that emphasize surface-level features (e.g., text-focused tasks) exhibit greater vulnerability than knowledge-intensive tasks, which require deeper semantic understanding and reasoning.
+Comparison of Judge Protocols.
 In this section, we compare two evaluation protocols used on LLM-as-a-Judge systems: pointwise scoring and pairwise comparison.
-The corresponding results are shown in Table III and Figure 5.
+The corresponding results are shown in Table III.
 Our findings show that while many adversarial strategies achieve high ASR under pointwise scoring, their P-ASR drops considerably in the pairwise comparison setting.
@@
 This discrepancy can be attributed to the nature of the pairwise protocol, which provides the judge model with an additional reference response for comparison.
-This allows the model to contrast the manipulated response with a stable reference, thereby reducing the attack’s impact.
 Despite the absolute differences in ASR and P-ASR, the relative effectiveness of attacks remains consistent for both protocols.
 Combined Attack (H6) and Fake Reasoning (H5) consistently rank among the most effective attacks in both ASR and P-ASR against all judge models, as shown in Table III.
-Conversely, optimization-based attack GCG (O5) and Auto-DAN (O6) exhibit limited impact on Openchat-3.5 (e.g., 0% and 6.67% ASR, respectively) under pointwise protocol, and correspondingly low P-ASRs (e.g., P-ASR <10% in most cases) under the pairwise protocol.
+Conversely, optimization-based attack GCG (O5) and Auto-DAN (O6) exhibit limited impact on Openchat-3.5 (e.g., 0% and 6.67% ASR, respectively) under the pointwise protocol, and correspondingly low P-ASRs (e.g., P-ASR <10% in most cases) under the pairwise protocol.
 $\bullet$ Finding 3: The relative effectiveness of attacks remains consistent across pointwise and pairwise protocols, i.e., attacks that achieve lower (or higher) ASR in one protocol also tend to yield lower (or higher) ASR in another.
-TABLE V: Evaluation of prevention-based defense methods (D1-D4) against adversarial attacks.
+TABLE V: The results of three detection-based defense methods (PPL, Windows PPL, and Naive LLM) evaluated against handcrafted attacks (H1–H8) and optimization-based attacks (O1–O7).
+The reported metrics include accuracy (ACC, %), recall (%), precision (%), and F1 score.
 Defense Metric Attack Methods Average
 H1 H2 H3 H4 H5 H6 H7 H8 O1 O2 O3 O4 O5 O6 O7
-D1 ASR(%) 0.000 0.000 0.000 0.000 3.330 100.00 90.00 3.330 3.330 3.330 3.330 23.33 6.670 6.670 6.670 16.67
-SDR -0.055 -0.084 -0.053 -0.126 -0.018 0.845 0.629 0.171 -0.041 -0.023 -0.031 0.054 -0.100 -0.036 -0.021 0.074
-iSDR -0.642 -0.698 -0.646 -0.574 -0.633 0.281 0.465 -0.336 -0.508 -0.417 -0.386 -0.150 -0.441 -0.405 -0.354 -0.363
-D2 ASR(%) 73.33 80.00 80.00 53.33 90.00 96.67 93.33 40.00 76.67 76.67 80.00 33.33 6.670 10.00 20.00 60.67
-SDR -0.070 -0.020 -0.016 -0.317 0.097 0.117 -0.108 -0.289 -0.049 -0.082 -0.126 -0.594 -0.634 -0.634 -0.517 -0.216
-iSDR 0.066 0.088 0.112 -0.043 0.204 0.275 0.450 -0.074 0.207 0.245 0.241 -0.075 -0.253 -0.281 -0.128 0.069
-D3 ASR(%) 76.67 86.67 86.67 46.67 96.67 100.0 86.67 50.00 83.33 83.33 90.00 33.33 0.000 6.670 23.33 63.33
-SDR -0.023 0.012 0.048 -0.329 0.106 0.147 -0.064 -0.215 0.001 -0.058 -0.040 -0.594 -0.768 -0.655 -0.533 -0.198
-iSDR 0.102 0.110 0.166 -0.065 0.203 0.295 0.485 -0.010 0.246 0.260 0.318 -0.085 -0.397 -0.312 -0.154 0.077
-D4 ASR(%) 80.00 80.00 80.00 46.67 93.33 100.0 90.00 43.33 80.00 76.67 73.33 26.67 13.33 10.00 26.67 61.33
-SDR -0.035 -0.033 -0.043 -0.329 0.045 0.119 -0.124 -0.287 -0.051 -0.126 -0.161 -0.662 -0.690 -0.713 -0.588 -0.245
-iSDR 0.114 0.089 0.099 -0.042 0.166 0.291 0.449 -0.058 0.218 0.216 0.220 -0.130 -0.296 -0.347 -0.185 0.054
-TABLE VI: Evaluation of detection-based defense methods (D5–D7) against adversarial attacks.
-Defense Metric Attack Methods Average
-H1 H2 H3 H4 H5 H6 H7 H8 O1 O2 O3 O4 O5 O6 O7
-D5 ACC(%) 55.00 53.33 51.67 50.00 48.33 50.00 71.67 23.64 46.67 31.67 28.33 51.67 51.67 55.00 56.67 48.35
+PPL ACC(%) 55.00 53.33 51.67 50.00 48.33 50.00 71.67 23.64 46.67 31.67 28.33 51.67 51.67 55.00 56.67 48.35
 Recall(%) 66.67 63.33 60.00 56.67 53.33 56.67 100.00 0.00 50.00 20.00 13.33 60.00 60.00 66.67 70.00 53.11
@@
 F1 0.597 0.576 0.554 0.531 0.508 0.531 0.779 0.000 0.484 0.226 0.157 0.554 0.554 0.597 0.618 0.484
-D6 ACC(%) 53.33 55.00 55.00 50.00 46.67 48.33 70.00 20.00 46.67 45.00 50.00 50.00 70.00 53.33 55.00 51.22
+Windows PPL ACC(%) 53.33 55.00 55.00 50.00 46.67 48.33 70.00 20.00 46.67 45.00 50.00 50.00 70.00 53.33 55.00 51.22
 Recall(%) 66.67 70.00 70.00 60.00 53.33 56.67 100.00 0.00 53.33 50.00 60.00 60.00 100.00 66.67 70.00 62.44
@@
 F1 0.588 0.609 0.609 0.545 0.500 0.523 0.769 0.000 0.500 0.476 0.545 0.545 0.769 0.588 0.609 0.545
-D7 ACC(%) 61.67 60.00 70.00 71.67 50.00 61.67 75.00 58.33 40.00 58.33 63.33 81.67 83.33 73.33 76.67 65.67
+Naive LLM ACC(%) 61.67 60.00 70.00 71.67 50.00 61.67 75.00 58.33 40.00 58.33 63.33 81.67 83.33 73.33 76.67 65.67
 Recall(%) 56.67 53.33 73.33 76.67 33.33 56.67 83.33 50.00 13.33 50.00 60.00 96.67 100.00 80.00 86.67 64.71
@@
 F1 0.597 0.571 0.710 0.730 0.400 0.596 0.769 0.545 0.182 0.545 0.621 0.841 0.857 0.750 0.788 0.627
-5.1.2 Defense Performance
-We evaluate 7 representative defenses methods from both prevention-based and detection-based paradigms.
-These include: retokenization[17] (D1), delimiter insertion (D2), sandwich prompting (D3), instruction augmentation (D4), perplexity filtering (PPL) (D5), windowed PPL (WinPPL) (D6), and a naive LLM-based detector (D7).
-Since prevention-based and detection-based methods are based on different mechanism and evaluated on different metrics, we report their results separately in Table V and Table VI, respectively.
-Prevention-based Defenses (D1–D4).
+TABLE VI: The results of four prevention-based defense methods (Retokenizer, Delimiters, Sandwich Prevention, and Instructional Prevention) evaluated against handcrafted attacks (H1–H8) and optimization-based attacks (O1–O7).
+The reported metrics include attack success rate (ASR, %), score difference rate (SDR), and improved score difference rate (iSDR).
+Defense Metric Attack Methods Average
+H1 H2 H3 H4 H5 H6 H7 H8 O1 O2 O3 O4 O5 O6 O7
+Retokenizer ASR(%) 0.000 0.000 0.000 0.000 3.330 100.00 90.00 3.330 3.330 3.330 3.330 23.33 6.670 6.670 6.670 16.67
+SDR -0.055 -0.084 -0.053 -0.126 -0.018 0.845 0.629 0.171 -0.041 -0.023 -0.031 0.054 -0.100 -0.036 -0.021 0.074
+iSDR -0.642 -0.698 -0.646 -0.574 -0.633 0.281 0.465 -0.336 -0.508 -0.417 -0.386 -0.150 -0.441 -0.405 -0.354 -0.363
+Delimiters ASR(%) 73.33 80.00 80.00 53.33 90.00 96.67 93.33 40.00 76.67 76.67 80.00 33.33 6.670 10.00 20.00 60.67
+SDR -0.070 -0.020 -0.016 -0.317 0.097 0.117 -0.108 -0.289 -0.049 -0.082 -0.126 -0.594 -0.634 -0.634 -0.517 -0.216
+iSDR 0.066 0.088 0.112 -0.043 0.204 0.275 0.450 -0.074 0.207 0.245 0.241 -0.075 -0.253 -0.281 -0.128 0.069
+Sandwich. ASR(%) 76.67 86.67 86.67 46.67 96.67 100.0 86.67 50.00 83.33 83.33 90.00 33.33 0.000 6.670 23.33 63.33
+SDR -0.023 0.012 0.048 -0.329 0.106 0.147 -0.064 -0.215 0.001 -0.058 -0.040 -0.594 -0.768 -0.655 -0.533 -0.198
+iSDR 0.102 0.110 0.166 -0.065 0.203 0.295 0.485 -0.010 0.246 0.260 0.318 -0.085 -0.397 -0.312 -0.154 0.077
+Instruction. ASR(%) 80.00 80.00 80.00 46.67 93.33 100.0 90.00 43.33 80.00 76.67 73.33 26.67 13.33 10.00 26.67 61.33
+SDR -0.035 -0.033 -0.043 -0.329 0.045 0.119 -0.124 -0.287 -0.051 -0.126 -0.161 -0.662 -0.690 -0.713 -0.588 -0.245
+iSDR 0.114 0.089 0.099 -0.042 0.166 0.291 0.449 -0.058 0.218 0.216 0.220 -0.130 -0.296 -0.347 -0.185 0.054
+V-A2 Defense Performance
+We evaluate 7 representative defense methods from both prevention-based and detection-based paradigms.
+These include perplexity filtering (PPL), windowed PPL (WinPPL), a naive LLM-based detector and retokenizationbaseDefense , delimiter insertion, sandwich prompting, and instruction augmentation.
+Since prevention-based and detection-based methods are based on different mechanisms and evaluated on other metrics, we report their results separately in Table VI and Table V, respectively.
+Detection-based Defenses Among detection-based methods, perplexity-based filters (PPL and WinPPL) generally underperform.
+They achieve relatively low accuracy (48.35% and 51.22%), recall (53.11% and 62.44%), precision (44.98% and 48.66%), and F1 scores (0.484 and 0.545), indicating limited capability in distinguishing between adversarial and non-adversarial inputs.
+In contrast, the naive LLM-based Detector achieves the highest detection performance across all metrics, with an accuracy (65.67%), recall (64.71%), precision (65.49%), and F1 score (0.627), outperforming all other prevention-based defenses, particularly against complex heuristic-based manipulations.
+Its effectiveness stems from leveraging a powerful LLM to identify subtle adversarial patterns in responses.
+However, this approach incurs significant computational overhead, as it requires querying the LLM detector for each input response.
+Prevention-based Defenses.
 Overall, prevention-based methods demonstrate moderate effectiveness across most attacks.
 However, they remain largely ineffective against stronger adversarial attacks, such as Combined Attack (H6) and AdvEval (H7).
-Among these methods, Retokenization (D1) yields the best average defense performance, achieving a reduced ASR of 16.67% and an iSDR of -0.394, significantly outperforming the other methods (D2–D4), which exhibit ASRs exceeding 60% and positive iSDR values above 0.03.
-Despite this relative success, prior findings [17] indicate that Retokenization (D1) can degrade model performance on benign inputs.
-This degradation arises from altered tokenization patterns, potentially affecting how the model interprets the content and leading to inaccurate evaluations.
+Among these methods, Retokenization yields the best average defense performance, with an ASR reduced to 16.67% and an iSDR of -0.363, significantly outperforming the other methods, which exhibit ASRs exceeding 60% and positive iSDR values above 0.03.
+Despite this relative success, prior findings baseDefense indicate that Retokenization can degrade model performance on benign inputs.
+This degradation arises from altered tokenization patterns, which can potentially affect how the model interprets the content and lead to inaccurate evaluations.
 Thus, while Retokenization offers strong adversarial robustness, it introduces a trade-off between defense effectiveness and benign input performance.
-Detection-based Defenses (D5–D7) Among detection-based methods, perplexity-based filters (PPL and WinPPL) generally underperform.
-They achieve relatively low accuracy (48.35% and 51.22%), recall (53.11% and 62.44%), precision (45.01% and 48.66%), and F1 scores (0.518 and 0.545), indicating limited capability in distinguishing adversarial inputs.
-In contrast, the naive LLM-based detector achieves the highest detection performance across all metrics, with an accuracy (65.67%), recall (64.71%), precision (65.49%) and F1 score (62.68%), outperforming all other prevention-based defenses, particularly against complex heuristic-based manipulations.
-Its effectiveness stems from leveraging a powerful LLM to identify subtle adversarial patterns in responses.
-However, this approach incurs significant computational overhead, as it requires querying the LLM detector for each input response.
 $\bullet$ Finding 4: Retokenization and the Naive LLM-based Detector are two of the most effective defense methods against adversarial attacks.
 However, their effectiveness comes with trade-offs: Retokenization may reduce performance on benign inputs due to altered tokenization patterns, while the LLM-based Detector introduces substantial computational overhead.
-5.2 Impact of Prompts and Model Choices (RQ2)
-TABLE VII: Comparison of robustness across different judge prompt templates on Text Translation tasks(T1).
-Higher ASR (weaker robustness)
-Lower ASR (stronger robustness)
+V-B Impact of Prompts and Model Choices (RQ2)
+TABLE VII: Comparison of robustness across different judge prompt templates on the Text Translation task (T1).
+Orange represents a higher P-ASR, suggesting weaker robustness across different attack methods.
+Green represents a lower average P-ASR, indicating stronger robustness across the four judge models.
 Judge Model H4 H5 H6 Avg
@@
 (d) Optimized Prompt (ours)
-5.2.1 Impact of Judge Prompt Template
-We examine the impact of prompt templates on the robustness of LLM-as-a-Judge systems.
-Specifically, we evaluate three widely used prompt templates, e.g., Vanilla Prompt, Arena-Hard Prompt and Google Vertex Promp, across four judge models on the translation task under three adversarial attacks (H4–H6).
-In this evaluation, a higher ASR indicate weaker robustness.
-The results are shown in Table VII (a-c).
+TABLE VIII: Our optimized prompt.
+System Prompt:
+You are an expert evaluator. You need to perform a translation task ({##SOURCE_LAN##} to {##TARGET_LAN##}).
+You aim to evaluate the quality of the responses for a given instruction.
+Your goal is to select the best response for the given instruction.
+Select Response A or Response B for the given instruction.
+Two different AI chatbots, respectively, generate the two outputs.
+Begin your evaluation by generating your own response to the prompt.
+You must provide your response before judging any responses.
+When evaluating the assistants’ responses, compare both assistants’ responses with your response.
+You must identify and correct any errors or inaccuracies.
+Please consider the helpfulness, relevance, accuracy, and level of detail of their responses.
+Please provide your explanation before providing your preference.
+User Prompt:
+### Instruction: {##SOURCE##}
+### Response A: {##A##}
+### Response B: {##B##}
+V-B1 Impact of Judge Prompt Template
+We investigate how different prompt templates affect the robustness of LLM-as-a-Judge systems.
+Specifically, we employ a pairwise evaluation protocol to compare three widely used templates—Vanilla Prompt, Arena-Hard Prompt, and Google Vertex Prompt—across four judge models in the Text Translation task (T1) under three adversarial attacks (H4–H6).
+In this evaluation, a higher P-ASR reflects weaker robustness.
+The results in Table VII (a)-(c) are averaged across six language pairs: Chinese–English, German–English, Yoruba–English, English–German, English–Yoruba, and English–Chinese.
 We observe that LLM-as-a-Judge systems are highly sensitive to the choice of prompt template.
-The same judge model can exhibit significantly different ASR values depending on the template used.
-For example, under H4, Mistral-7B yields an ASR of 37.50% with the Vanilla prompt, which increases to 96.25% with the Arena-Hard prompt.
-Similarly, LLama-3.1-8B displays a wide range in ASR, from 6.67% ASR (Vanilla) to 73.33% (Google Vertex) under H4, further highlighting this sensitivity.
-To improve robustness, we apply the coordinate ascent strategy described in Algorithm 1 to derive an optimized prompt.
-Initializing from the Arena-Hard template, we iteratively refine individual prompt components based on ASR under the H4 attack.
-The optimization is performed on the translation task, with chinese2english language pair.
-We conduct three iterations of coordinate ascent using Openchat-3.5 as the target model.
-The optimized prompt is then evaluated for generalization across five additional language pairs within the translation task, and the transferability under H5 and H6 attacks.
+The same judge model can exhibit significantly different P-ASR values depending on the template used.
+For example, under H4, Mistral-7B yields a P-ASR of 37.50% with the Vanilla prompt, which increases to 96.25% with the Arena-Hard prompt.
+Similarly, LLama-3.1-8B exhibits a wide range in P-ASR, from 6.67% P-ASR (Vanilla) to 73.33% (Google Vertex) under H4, further highlighting this sensitivity.
+To enhance robustness, we employ the coordinate ascent strategy described in Algorithm 1 to derive an optimized prompt.
+Starting from the Arena-Hard template, individual prompt components are iteratively replaced based on their P-ASR under a relatively strong H4 attack.
+This optimization is performed on the Text Translation task (T1) using the Chinese–English language pair.
+We conduct three iterations of coordinate ascent with Openchat-3.5 as the target model.
+The resulting optimized prompt is subsequently evaluated for generalization across five additional language pairs (German–English, Yoruba–English, English–German, English–Yoruba, English–Chinese) within T1, as well as for transferability under the H5 and H6 attacks, which represent the other stronger attacks in the set.
 Compared with the baselines, our optimized prompt achieves superior robustness (Table VII (d)).
-On Openchat-3.5, it reduces the average ASR across H4–H6 to 6.11%, outperforming Arena-Hard (20.08%) and Google Vertex (23.06%), while also maintaining comparably low ASR on Qwen-2.5 and LLaMA-3.1-8B.
-Notably, for Mistral-7B, it brings the ASR under H4 down from 96.25% (Arena-Hard) to only 14.17%, demonstrating a substantial improvement in robustness.
-The optimized prompt template is presented in Table X (appendix).
-Importantly, unlike prior work that relies on extensive prompt search or model fine-tuning, our method performs optimization in a highly cost-efficient manner, targeting a single attack (H4), a single task (translation), and one language pair (chinese2english).
-Despite this constrained budget, the optimized prompt generalizes well on additional attacks (H5, H6) and models (Openchat, Qwen, Mistral, LLaMA), highlighting the strong generalizability and transferability of our approach.
+On Qwen2.5-7B, it reduces the average P-ASR across H4–H6 to 3.61%, outperforming Arena-Hard (15.42%) and Google Vertex (9.76%), while also maintaining a comparably low P-ASR on Mistral-7B and LLaMA-3.1-8B.
+Notably, for Mistral-7B, it brings the P-ASR under H4 down from 96.25% (Arena-Hard) to only 14.17%, demonstrating a substantial improvement in robustness.
+The optimized prompt template is presented in Table VIII.
+Importantly, unlike prior work that relies on extensive prompt search or model fine-tuning, our method performs optimization in a highly cost-efficient manner, targeting a single attack (H4), a single task (T1), and one language pair (Chinese-English).
+Despite this constrained budget, the optimized prompt generalizes well on additional attacks (H5, H6) and models, highlighting the strong generalizability and transferability of our approach.
 This exhibits that lightweight, targeted prompt modifications can yield broad defensive benefits and robustness for LLM-as-a-Judge systems.
-$\bullet$ Finding 5: The robustness of LLM-as-a-Judge systems is highly sensitive to the choice of prompt template.
-$\bullet$ Finding 6: We propose a low-cost yet highly transferable prompt template optimization method that improves robustness.
-The optimized template consistently outperforms existing prompt template.
+[…]
+[…]
 (a) Empty Attack
@@
 (h) Average
-Figure 3: Evaluate the robustness of different judge models on 7 attacks.
-5.2.2 Impact of Judge Model
+Figure 4: Robustness evaluation of different judge models against 7 adversarial attacks.
+Each subplot shows the Attack Success Rate (ASR), Success Defense Rate (SDR), and Improved Success Defense Rate (iSDR) for a specific attack method, with (h) presenting the average performance across all attacks.
+V-B2 Impact of Judge Model
 We evaluate a range of judge models, including GPT-4o, Openchat, JudgeLM-7/13B, Prometheus-7B, DeepSeek-R1, PandaLM-7B, and AutoJ-13B, comparing their robustness against various adversarial attacks.
-The results are summarized in Figure  3.
-GPT-4o, considered to be the defacto LLM-as-a-Judge system, demonstrates strong robustness across different adversarial attacks, achieving an ASR of 71.26% and an iSDR of 0.1235 (Figure 3h).
+The results are summarized in Figure 4.
+GPT-4o, considered to be the extensive LLM-as-a-Judge system, demonstrates strong robustness across different adversarial attacks, achieving an ASR of 71.26% and an iSDR of 0.1235 (Figure 3h).
 These results validate the current practice of using GPT-4o as a judge model and suggest that it offers a relatively reliable evaluation setup.
@@
 This highlights the effectiveness of task-specific fine-tuning in improving the robustness of LLM-as-a-Judge systems.
-Given its open-source availability and lower computational overhead compared to GPT-4o, JudgeLM-13B presents a promising and cost-effective alternative for reliable LLM-as-a-Judge system.
+Given its open-source availability and lower computational overhead compared to GPT-4o, JudgeLM-13B presents a promising and cost-effective alternative for a reliable LLM-as-a-Judge system.
 In contrast, reasoning-oriented models such as DeepSeek-R1 have been hypothesized to offer superior robustness in judgment tasks due to their enhanced reasoning capabilities.
 However, our results show that DeepSeek-R1 achieves only moderate robustness, with an ASR of 75.20% and an iSDR of 0.5643, indicating no significant advantage over other models in adversarial scenarios.
-$\bullet$ Finding 7: Judge-tuned models demonstrate the strongest robustness against adversarial attacks due to alignment-focused fine-tuning.
-$\bullet$ Finding 8: Reasoning-focused models (e.g., DeepSeek-R1) are more susceptible to severe scoring errors when attacks succeed, indicating weaker reliability under adversarial conditions.
-5.3 Real-World Case Study (RQ3)
-We conduct experiments on a real-world industrial LLM-as-a-Judge platform, i.e., PAI-Judge [^2] https://pai.console.aliyun.com/#/ai-service/judge/overview.
-This platform provides an overall evaluation score for each input, along with a set of subscores reflecting multiple assessment dimensions, such as accuracy, fluency, and consistency.
-According to its official documentation, PAI-Judge also offers a premium version PAI-Judge-Plus, which is built on a larger model, and is reported to deliver superior judgment performance.
-Our initial evaluation targets PAI-Judge using pointwise protocol and includes both heuristic-Based attacks (e.g., Naive Attack, Escape Characters Attack, Context Ignore Attack, Fake Completion Attack, Fake Reasoning, Combined Attack (H1-H8)), and optimization-based attacks (PAIR (O2)).
-To ensure compatibility with the platform, we carefully adapt the prompt template to match the PAI-Judge’s format.
-However, our results show that these conventional adversarial strategies had minimal effect on the platform’s evaluation outcome.
-This suggests that PAI-Judge incorporates effective internal defense mechanisms, making it relatively robust to standard adversarial manipulations.
-Sample evaluation results are presented in Table VIII (appendix).
+$\bullet$ Finding 7: Judge-tuned models demonstrate the strongest robustness against adversarial attacks.
+In particular, JudgeLM-13B consistently achieves robustness levels that not only surpass other open-source judges but also exceed GPT-4o, highlighting the effectiveness of alignment-focused fine-tuning for enhancing adversarial resistance.
+V-C Real-World Case Study (RQ3)
+To demonstrate the practical utility of RobustJudge in identifying vulnerabilities in LLM-as-a-Judge systems, we conduct a comprehensive evaluation on PAI-Judge [^5] https://pai.console.aliyun.com/#/ai-service/judge/overview, a commercial industrial judging platform by Alibaba Cloud.
+PAI-Judge provides multi-dimensional evaluation through a hierarchical scoring mechanism, generating individual subscores (0-10 scale) for accuracy, fluency, consistency, relevance, and helpfulness, which are aggregated into an overall score via a weighted combination.
+The platform offers two variants: the standard version and PAI-Judge-Plus.
+Our evaluation follows a systematic three-phase approach.
+We first deploy conventional adversarial attacks under the pointwise protocol, including heuristic-based attacks (Naive Attack, Escape Characters Attack, Context Ignore Attack, Fake Completion Attack, Fake Reasoning) and optimization-based attacks (PAIR (O2)).
+We adapt prompt templates to match PAI-Judge’s API format.
+However, these standard strategies have a minimal impact—both subscores and overall scores remain largely unchanged across all tested variants, suggesting that PAI-Judge incorporates effective internal defenses, such as prompt filtering, output sanitization, or adversarial training.
+We leverage our RobustJudge framework with gradient-based optimization, employing Qwen-2.5-7B-Instruct fine-tuned using GCG-like methods to craft adversarial responses that appear plausible while containing quality flaws.
+The optimization minimizes $\mathcal{L}_{\text{adv}}=-\log P_{\text{judge}}(\text{high score}\mid\text{query},\text{adversarial response})$, yielding samples that exploit the judge’s decision boundaries through semantic manipulations, fabricated reasoning, and misleading cues.
+We then introduce long-suffix augmentation, which involves appending lengthy, irrelevant content (such as repetitive filler text, irrelevant facts, and distracting narratives) to exploit the judge’s attention mechanism limitations.
+By overwhelming the model with excessive input length, we hypothesize that the judge fails to focus on core adversarial content, inflating scores.
+Our results reveal a critical vulnerability in both PAI-Judge platforms (Figure 5).
+The composite attack—combining Qwen-2.5-7B-optimized adversarial content with long-suffix augmentation—achieves remarkably high scores despite responses containing significant flaws warranting near-zero scores.
+Standard PAI-Judge exhibits severe susceptibility: with suffix lengths of 1000 characters, average subscore increases from approximately 1.5 (baseline) to over 6.0, with overall score reaching 7.2.
+PAI-Judge-Plus maintains low scores below 1500 characters but succumbs once suffix length exceeds 2000 characters, achieving an average subscore of 5.8 and an overall score of 6.5, indicating acceptable quality in the platform’s rubric.
 (a) PAI-Judge
 (b) PAI-Judge-Plus
-Figure 4: Comparison of Avg Subscore and Overall Score for PAI-Judge and PAI-Judge-Plus.
-To further probe the system’s robustness, we extended the PAIR-optimized adversarial prompts by appending long-suffix manipulations (lengthy irrelevant or distracting content).
-Surprisingly, this composite attack results in high scores on both PAI-Judge (a) and PAI-Judge-Plus (b), despite that these adversarial responses should have received near-zero scores (Figure 4).
-We further investigate the attack by gradually increasing the length of the appended suffix and observe a rising trend in effectiveness.
-While PAI-Judge-Plus demonstrates some resistance, its scoring eventually increases significantly once the suffix length approaches 2000 letters.
-These results reveal a critical loophole in PAI-Judge and PAI-Judge-Plus platform.
-Using our RobustJudge framework, we are able to systematically evaluate and uncover this vulnerability in a real-world judge system.
-These findings demonstrate the practical utility of RobustJudge in identifying hidden flaws and guiding the development of more robust LLM-as-a-Judge systems.
-We responsibly disclosed this vulnerability to the PAI-Judge platform team prior to the release of our results.
-$\bullet$ Finding 9: Conventional adversarial attacks were largely ineffective against the industrial PAI-Judge platform.
-$\bullet$ Finding 10: We identify a loophole in both PAI-Judge and PAI-Judge-Plus, where combining PAIR-optimized adversarial inputs with long-suffix can compromise the platform’s judgment reliability.
-6 Conclusion
-This work presents the first scalable and fully automated framework to evaluate the robustness and reliability of LLM-as-a-Judge systems across multiple attack scenarios.
-We systematically benchmarked state-of-the-art LLM-based evaluators under various adversarial settings and found that they are vulnerable to manipulation, often producing biased or incorrect judgments when exposed to crafted inputs.
-Based on these findings, we conclude that current LLM-as-a-Judge systems are not yet sufficiently robust for secure deployment in critical evaluation tasks, and highlight the need for future research to strengthen their resilience against adversarial attacks.
-Our framework and benchmarks provide a valuable tool for the community to assess and track the progress of LLM-based evaluators in adversarial robustness.
+Figure 5: Impact of attack on PAI-Judge variants.
+Progressive length analysis (0-3000 characters, 200-character increments) reveals sharp transition points where scores increase rapidly—around 800-1000 characters for PAI-Judge and 1800-2000 characters for PAI-Judge-Plus (2× better resistance).
+Beyond these thresholds, scores plateau at 7-8 and 6-7, respectively.
+Attack effectiveness varies across dimensions: fluency and consistency subscores increase more dramatically (7-8) than accuracy scores (5-6), indicating attention degradation primarily affects surface-level assessments rather than semantic understanding.
+$\bullet$ Finding 8: Conventional adversarial attacks were largely ineffective against the industrial PAI-Judge platform.
+$\bullet$ Finding 9: We identify a loophole in both PAI-Judge and PAI-Judge-Plus, where combining PAIR-optimized adversarial inputs with long suffixes can compromise the platform’s judgment reliability.
+VI Conclusion
+This work presents the first scalable, automated framework for evaluating LLM-as-a-Judge robustness under adversarial attacks.
+Our benchmarking shows that current evaluators (even Alibaba’s PAI platform) remain vulnerable and biased, underscoring the need for stronger defenses and more reliable real-world deployment.
+We also discovered that the robustness varies substantially across prompt templates.
+Our findings provide actionable insights for deploying trustworthy LLM-as-a-Judge systems in security-sensitive applications.